### v1

# NN-ResDMD: Learning Koopman Representations for Complex Dynamics with Spectral Residuals

Abstract
Analyzing long-term behaviors in high-dimensional nonlinear dynamical systems remains a significant challenge. The Koopman operator framework has emerged as a powerful tool to address this issue by providing a globally linear perspective on nonlinear dynamics. However, existing methods for approximating the Koopman operator and its spectral components, particularly in large-scale systems, often lack robust theoretical guarantees. Residual Dynamic Mode Decomposition (ResDMD) introduces a spectral residual measure to assess the convergence of the estimated Koopman spectrum, which helps filter out spurious spectral components. Nevertheless, it depends on pre-computed spectra, thereby inheriting their inaccuracies. To overcome its limitations, we introduce the Neural Network-ResDMD (NN-ResDMD), a method that directly estimates Koopman spectral components by minimizing the spectral residual. By leveraging neural networks, NN-ResDMD automatically identifies the optimal basis functions of the Koopman invariant subspace, eliminating the need for manual selection and improving the reliability of the analysis. Experiments on physical and biological systems demonstrate that NN-ResDMD significantly improves both accuracy and scalability, making it an effective tool for analyzing complex dynamical systems.
1 Introduction
In the study of complex dynamical systems, a critical challenge lies in accurately extracting and analyzing long-term behavior in high-dimensional nonlinear systems. Various data-driven methods [1, 2, 3, 4, 5, 6] have been developed to address this challenge, with the Koopman operator [7, 8] framework emerging as a powerful tool due to its ability to globally linearize nonlinear systems. Unlike local linearization methods [9, 10], which approximate dynamics near fixed points, the Koopman operator transforms the entire system into a linear form within an infinite-dimensional space, which allows the use of spectral analysis techniques to study complex dynamics.
Despite its promise, practical computational challenges arise from the infinite-dimensional nature of the Koopman operator. Numerical methods such as Extended Dynamic Mode Decomposition (EDMD) [11] have been developed to approximate the Koopman operator using a finite set of observables, making it possible to extract dynamic modes from data. However, EDMD lacks theoretical guarantees of convergence and may fail to capture the full Koopman spectrum accurately, particularly in large-scale, complex systems.
To address these limitations, Residual Dynamic Mode Decomposition (ResDMD) [12] was introduced, offering convergence guarantees through a spectral residual measure that quantifies how well the estimated Koopman spectrum converges to the true spectrum. By assessing convergence, ResDMD eliminates spurious spectral components that do not represent the system’s true dynamics, enhancing the reliability of spectral estimation. However, ResDMD primarily filters precomputed spectra rather than directly approximating Koopman spectra, limiting its ability to independently refine spectral estimates.
In this paper, we propose Neural Network-ResDMD (NN-ResDMD), which overcomes this limitation by providing a method to directly compute Koopman eigenpairs by minimizing the spectral residual. Additionally, NN-ResDMD employs neural networks to automatically select basis functions, eliminating the need for manual intervention, a common challenge in EDMD-based methods. Through experiments on both toy models and real-world high-dimensional systems, we demonstrate that NN-ResDMD significantly improves accuracy and scalability, making it a practical and effective tool for analyzing complex dynamical systems.
2 Preliminary on Koopman operator
Consider a discrete-time dynamical system $(\Omega,\mu)$ governed by a map $F:\Omega\to\Omega$, where $\Omega\subseteq\mathbb{R}^{d}$ is the state space, and $\mu$ is a probability measure. The evolution of the system is described by:
$x_{k+1}=F(x_{k}),\quad k\in\mathbb{Z}^{+}.$
The Koopman operator $\mathcal{K}$ acts on observables $g\in L^{2}(\Omega,\mu)$ as:
$\mathcal{K}g=g\circ F.$
Although $F$ is nonlinear, the Koopman operator $\mathcal{K}$ is linear, enabling spectral analysis of the system in the infinite-dimension function space.
A key aspect of modern Koopman operator theory is Koopman Mode Decomposition (KMD) [6], which represents system dynamics through its spectral components, i.e. the eigenvalues, Koopman modes, and eigenfunctions. The discrete spectrum is particularly important for insights into long-term behavior, such as periodicity and stability. Our analysis emphasizes these spectral components derived from KMD. Specifically, we seek eigenpairs $(\lambda_{i},\phi_{i})$, where $\lambda_{i}$ are eigenvalues and $\phi_{i}$ are the corresponding Koopman eigenfunctions.
One of the most prominent numerical methods to approximate the Koopman operator and its spectral components is the Extended Dynamic Mode Decomposition (EDMD) method, introduced by [11]. In EDMD, a set of observables (dictionary or basis functions) $\mathbf{\Psi}=[\psi_{1},\dots,\psi_{N_{K}}]$ is selected, and the span of these observables defines the subspace $V_{N_{K}}\coloneqq\text{span}\{\psi_{i}\}_{i=1}^{N_{K}}$. Snapshots of the system’s state are then collected, and the method constructs a finite-dimensional approximation of the Koopman operator by solving a least-squares problem that relates the snapshots of observables. This enables the computation of eigenvalues, eigenfunctions, and Koopman modes. Note that while common choices of dictionary functions are polynomials, Fourier basis, RBF functions, etc., the optimal choice of basis functions is usually unknown a priori and depends heavily on the specific dynamical system.
Given independent and identically distributed data snapshots $\{(x_{i},y_{i})\}_{i=1}^{m}$ with $y_{i}=F(x_{i})$, two matrices $\Psi_{X}$ and $\Psi_{Y}$ are formed by evaluating the dictionary on the data snapshots:
$\Psi_{X}=\begin{bmatrix}\psi_{1}(x_{1})&\dots&\psi_{N_{K}}(x_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(x_{m})&\dots&\psi_{N_{K}}(x_{m})\end{bmatrix},\quad\Psi_{Y}=\begin{bmatrix}\psi_{1}(y_{1})&\dots&\psi_{N_{K}}(y_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(y_{m})&\dots&\psi_{N_{K}}(y_{m})\end{bmatrix}.$
EDMD computes the Koopman matrix approximation as $K=\Psi_{X}^{\dagger}\Psi_{Y}$, where $\Psi_{X}^{\dagger}$ is the pseudo-inverse of $\Psi_{X}$. The eigenvalues of $K$ provide approximations of the Koopman operator’s spectrum, and the Koopman eigenfunctions $\phi_{i}$ are approximated as $\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i}$, where $\mathbf{v}_{i}\in\mathbb{C}^{N_{K}}$ is the $i$-th eigenvector of $K$.
3 Koopman Operator Learning
While EDMD approximates the Koopman operator, it suffers from spectral pollution as increasing dictionary sizes introduce spurious eigenvalues, obscuring the system’s true dynamics. Residual Dynamic Mode Decomposition (ResDMD) [12] addresses this by filtering out spurious eigenvalues using spectral residuals but relies on precomputed eigenpairs, inheriting inaccuracies from methods like EDMD.
To overcome these limitations, we propose Neural Network-ResDMD (NN-ResDMD), which directly approximates the Koopman operator and its spectrum by minimizing a ResDMD-specific loss function. Using a Feedforward Neural Network (FNN), NN-ResDMD also optimizes dictionary functions for the Koopman invariant subspace, eliminating the need for manual basis selection.
3.1 ResDMD Review
Now, suppose we have obtained an eigenpair $(\lambda,\phi)$ of $\mathcal{K}$ from EDMD or other methods [13, 14, 15, 16, 17, 18] where $\lambda\in\mathbb{C}$ and the eigenfunction $\phi$ is expanded in terms of dictionary functions, i.e., $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}v_{i}\in V_{N_{K}}$ for some $\mathbf{v}\in\mathbb{C}^{N_{K}}$, where $v_{i}$ represents weights of the span. Without loss of generality, we consider $\phi$ has been normalized, i.e., $\|\phi\|_{2}=1$. The accuracy of this eigenpair approximation in the ResDMD framework can be measured by computing its squared relative residual using the dictionary in the following way:
[…]
$=\sum_{i,j=1}^{N_{K}}\bar{v}_{i}\left[\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\lambda\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\bar{\lambda}\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}+|\lambda|^{2}\langle\psi_{i},\psi_{j}\rangle_{\mu}\right]v_{j},$ (3.1)
where $\bar{v}_{i},\bar{\lambda}$ denote the complex conjugate of $v_{i},\lambda$.
This squared relative residual in (A.1) is the theoretical value that measures the distance between $\phi$ and the eigenspace associated with $\lambda$, especially under the assumption that $\lambda$ is in the discrete spectrum of $\mathcal{K}$. To approximate this residual in practice, we apply the Galerkin approximation [19], which states that as the number of data points $m$ increases, the following limits hold:
$\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{X}^{*}\Psi_{X}\right]_{ij}$ $=\langle\psi_{i},\psi_{j}\rangle_{\mu},$ (3.2)
$\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{X}^{*}\Psi_{Y}\right]_{ij}$ $=\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu},$
$\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{Y}^{*}\Psi_{Y}\right]_{ij}$ $=\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}=\langle\psi_{i},\mathcal{K}^{*}\mathcal{K}\psi_{j}\rangle_{\mu},$
where * denotes complex conjugate. Using this approximation, the squared relative residual from (A.1) is approximated as follows (see A.1 for more details):
$\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}\left[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}-\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}\right]\mathbf{v}.$ (3.3)
where (3.3), denoted as $\widehat{\textit{res}}(\lambda,\phi)^{2}$, represents the approximation of the theoretical value in (A.1).
By definition in (A.1), the residual quantifies the deviation from the spectral property, measuring how far the estimated eigenpair is from the true spectrum. In practice, (3.3) is calculated for all precomputed eigenpairs, retaining those with residuals below a threshold. However, while residuals help filter and select valid eigenpairs, they do not improve the accuracy of eigenpair estimation.
3.2 Neural Network-ResDMD
General framework In this section, we present the Neural Network-ResDMD (NN-ResDMD) framework, designed to compute the eigenpairs of the Koopman operator directly using ResDMD-based spectral residuals, as illustrated in Figure 1. The method first determines the optimal dictionary functions by minimizing the total residual $J\coloneqq\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}$, over all computed eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{N_{K}}$. The spectral residual directly impacts the finite-dimensional projection of the Koopman operator and our method minimizes this residual to ensure the learned basis functions adequately capture the Koopman dynamics. This approach allows the construction of the Koopman operator matrix $\tilde{K}$ without relying on external methods or post-processing. Equation (3.5) enables NN-ResDMD to compute eigenpairs directly, improving accuracy compared to ResDMD, which relies on filtering precomputed results from other methods.
In this framework, neural networks parameterize the dictionary functions $\mathbf{\Psi}(x;\theta)$, where $\theta$ represents the network parameters. By minimizing the spectral residual $J$, this approach directly optimizes the dictionary functions towards better approximation of Koopman spectral components, which ensures the learned operator captures the underlying spectral properties of the dynamical system. This is fundamentally different from traditional methods like EDMD, which focus on minimizing prediction errors in the observable space without explicitly considering spectral accuracy. The neural network architecture serves as a flexible function approximator, that allows the framework to adaptively learn the optimal dictionary that minimizes spectral residuals, thereby producing more accurate and reliable Koopman spectral decompositions. This spectral-oriented optimization improves the accuracy of eigenvalues approximations and enhances the quality of the computed eigenfunctions, which leads to better characterization of the system’s dynamic behavior.
Figure 1: (Left) The classical ResDMD and (Right) the Neural Networks based ResDMD methods
From Residual to NN This section explains how neural networks are integrated into the ResDMD framework. In ResDMD, the squared relative residual approximation (3.3) measures how well a computed eigenpair fits the dataset. If the Koopman matrix $K$ is well-approximated by the projected Koopman operator $\mathcal{K}_{N_{K}}$, the total residual $J$ should approach zero as more data is provided. Thus, $J$ can be used as a loss function, and the optimal Koopman matrix $\tilde{K}$ is obtained by minimizing:
$J=\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}.$ (3.4)
which is equivalent to minimization the following (See A.2 for more details):
$J=\frac{1}{m}\|(\Psi_{Y}-\Psi_{X}K)V\|_{F}^{2}$ (3.5)
where $V$ is a matrix in which each column is an eigenvector $\mathbf{v}_{i}$ of Koopman matrix $K$. Thus, with a fixed dictionary function $\mathbf{\Psi}$, the explicit form for the optimal Koopman matrix $\tilde{K}$ can be directly computed as
$\tilde{K}=G^{\dagger}A$ (3.6)
where $G=\frac{1}{m}\Psi_{X}^{*}\Psi_{X},A=\frac{1}{m}\Psi_{X}^{*}\Psi_{Y}$.
Remark.
Typically a regularization term is needed to enhance stability. Here we add a small perturbation, i.e., $\tilde{K}=(G+\sigma I)^{-1}A$ for some small number $\sigma>0$.
As shown in (3.6), NN-ResDMD provides an explicit expression for $\tilde{K}$ given the optimal dictionary function $\mathbf{\Psi}$, allowing for the direct computation of Koopman eigenpairs. The optimization problem in Equation 3.5 is to minimize the error along the eigen-basis, in contrast to the optimization problem $\|\Psi_{Y}-\Psi_{X}K\|_{F}^{2}$ for EDMD, thereby yielding different optimal $\mathbf{\Psi}$ compared to EDMD. Therefore, although the $K$ update procedure appear identical to the EDMD approach, they originate from different theoretical foundations and serve different optimization purposes. Additionally, it automatically optimizes basis functions using neural networks, removing the need for manual selection. Since NN-ResDMD is based on the ResDMD framework, it also retains the theoretical convergence guarantees that EDMD lacks: EDMD has convergence results under strong assumptions, such as requiring the Koopman operator to be bounded (Assumption 2 in [20]), ResDMD requires only that the operator is closed and densely defined.
In NN-ResDMD, neural networks parameterize the dictionary functions $\mathbf{\Psi}(x;\theta)$ to minimize the total residual $J(\theta)$, as defined in (3.4). The feedforward neural network generates the dictionary functions based on data snapshots, and the total residual is given by:
$J(\theta)=\frac{1}{m}\|(\Psi_{Y}(\theta)-\Psi_{X}(\theta)K(\theta))V(\theta)\|_{F}^{2}$ (3.7)
where $K(\theta)$ and $V(\theta)$ depend on $\theta$. The Koopman matrix $\tilde{K}(\theta)$ is computed as:
$\tilde{K}(\theta)=G(\theta)^{\dagger}A(\theta)$ (3.8)
with $G(\theta)=\frac{1}{m}\Psi_{X}(\theta)^{*}\Psi_{X}(\theta)$ and $A(\theta)=\frac{1}{m}\Psi_{X}(\theta)^{*}\Psi_{Y}(\theta)$.
The algorithm alternates between updating $K(\theta)$ via least squares and optimizing $\theta$ using gradient descent until $J(\theta)$ converges, yielding the approximated Koopman spectrum and optimized dictionary functions. While it is possible to optimize both $K(\theta)$ and $J(\theta)$ simultaneously, as done in [21] and [22], our separate procedure ensures computational efficiency and numerical stability compared to the coupled optimization case.
Computing Algorithm In our neural networks implementation, we include some non-trainable basis outputs to enhance the dictionary functions. Specifically, we add a vector of ones and the coordinates of the state space as non-trainable basis in the output layer, which help avoid trivial solutions, i.e., $J=0$ for some initial $\theta$. For the network architecture, we build a three-layer Feedforward Network where each hidden layer size can be specified during training. We use the hyperbolic tangent (tanh) function as the activation function for the hidden layers. In terms of optimization, we employ the Adam optimizer for updating the network parameters. Adam is particularly well-suited for this task due to its ability to adapt the learning rate for each parameter, which can lead to faster convergence in the alternating optimization process between the network parameters and the Koopman matrix. The computing steps are illustrated in the following Algorithm 1.
Input: Dataset $X,Y$, number of observables $N_{K}$, learning step $\delta$, regularization parameter $\sigma$, loss function threshold $\epsilon>0$, grid points $\{z_{1},\ldots z_{n_{z}}\}$.
1 Initialize $\theta$, thus initializing $\mathbf{\Psi}(\theta)$ ;
2 Compute $\tilde{K}(\theta)$ and its eigenvector matrix $V(\theta)$ ;
3 while $J(\theta)>\epsilon$ do
4       Update $\theta=\theta-\delta\nabla_{\theta}J(\theta)$ ;
5       Compute $G(\theta)=\frac{1}{m}\Psi_{X}^{*}\Psi_{X},A(\theta)=\frac{1}{m}\Psi_{X}^{*}\Psi_{Y}$;
6       Update $\tilde{K}(\theta)=(G(\theta)+\sigma I)^{-1}A(\theta)$ and $V(\theta)$ ;
7
8
Output: $\tilde{K}(\theta)$, eigenpairs $\{(\lambda_{i},\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i})\}_{i=1}^{N_{K}}$ and pseudospectrum $\{z_{j}:\tau_{j}<\varepsilon\}$.
9
Algorithm 1 NN-ResDMD
While the practical advantages of NN-ResDMD are demonstrated through experiments, it is important to note its computational demands. The algorithm’s computational complexity stems primarily from its iterative optimization process. Each iteration involves a gradient descent update with complexity scaling linearly with both system dimensionality and neural network parameters. Although individual gradient steps are computationally lightweight for standard network architectures, the algorithm’s efficiency issue lies in its repeated least-squares optimizations. Compared to standard single least-squares computation as in most numerical algorithms, NN-ResDMD requires multiple iterations to achieve convergence, with stochastic gradient descent methods showing a theoretical $O(1/n)$ convergence rate. However, the method’s nonlinear optimization nature also presents challenges for establishing concrete convergence bounds and error estimates.
If the continuous spectrum of the Koopman operator is of interest, following the ResDMD paper’s idea, we can scan candidate spectrum values within a grid in the complex plane using the residuals. Specifically, we compute $\tau_{j}=\min_{\mathbf{v}_{i}\in\mathbb{C}^{N_{k}}}\widehat{\textit{res}}(z_{j},\mathbf{\Psi}(\theta)\mathbf{v}_{i})$, where $\tau_{j}$ is the minimum residual for a grid point $z_{j}\in\mathbb{C}$. The approximated whole spectrum containing the continuous spectrum is then given by $\{z_{j}:\tau_{j}<\varepsilon\}$. More details can be found in [12].
While the practical advantages of NN-ResDMD are demonstrated through experiments, it’s also worth noting that the method has theoretical underpinnings [23, 24] that support its convergence properties. A brief discussion on the convergence aspects of NN-ResDMD, leveraging existing results from approximation theory in Barron spaces, is provided in Appendix A.3. This discussion offers insights into how the neural network component of NN-ResDMD contributes to its effectiveness in approximating complex dynamical systems.
4 Application in physical and biological systems
In this chapter, we present three examples demonstrating the effectiveness of NN-ResDMD in estimating the key quantities of Koopman Mode Decomposition (KMD): spectrum, eigenfunctions, and Koopman modes. In the first low-dimensional example of a classical pendulum system, our method requires significantly fewer dictionary observables than [12, Section 4.3.1, Section 6.3] to compute the Koopman spectrum. The second high-dimensional example on turbulence highlights our method’s ability to detect acoustic vibrations and distinguish the pressure field through Koopman modes. The third example, a real-world high-dimensional neural system, compares NN-ResDMD with three popular methods—Hankel-DMD [25], EDMD with RBF basis, and kernelized-ResDMD [12]—and demonstrates its superiority in identifying and clustering latent dynamic structures. Together, these examples showcase NN-ResDMD’s performance across diverse systems and comprehensively evaluate its capabilities.
Specifically, in all three experiments, we compare NN-ResDMD with Hankel-DMD, a theoretically grounded method that analyzes dynamical systems using time-delayed state measurements (see Appendix A.6.1 for a justification). Although its performance rivals NN-ResDMD in the simple pendulum system, it fails to capture key dynamics in higher-dimensional systems.
4.1 Pendulum
The pendulum system is a measure-preserving system due to its Hamiltonian nature, which theoretically implies its whole spectrum lies on the unit circle. For its dynamical behaviors, if the initial position of the pendulum is sufficiently far from the peak and the initial angular speed sufficiently small, the pendulum will oscillate; otherwise, the pendulum will pass the peak and rotate. In other words, this complex system exhibits two types of dynamical behaviors: rotation and oscillation. Here we simulate two cases with different numbers of initial points. We choose 90 and 240 initial points uniformly in the domain $[-\pi,\pi]_{per}\times[-15,15]$. Each point evolves 1000 steps with a step size of 0.5. Thus, the total data size in each set is approximately $9\times 10^{4}$ and $2.4\times 10^{5}$, respectively.
As shown in Figure 4, only $N_{K}=300$ observables are needed to approximate the full spectrum, significantly fewer than the nearly 1000 required in [12, Section 4.3.1] for the same data size. Even with a much larger data size (Figure 4), the required observables remain small ($N_{K}=350$), demonstrating the robustness of efficient observables across data sizes.
As shown in Figure 4, we compare our method to four approaches: EDMD, EDMD with Dictionary Learning (EDMD-DL), Hankel-DMD, and ResDMD, on the dataset with 90 initial points to compute the Koopman matrix and its corresponding spectral information. The first three methods (EDMD [11], EDMD-DL [18], and Hankel-DMD are limited to computing eigenvalues associated with the point spectrum. In these experiments, both EDMD and ResDMD use the hyperbolic cross approximation with Hermite functions up to order 15 and Fourier functions up to order 20. Hankel-DMD uses a time delay of 150. While Hankel-DMD yields accurate eigenvalues and shows points near the unit circle which matches the ground truth for this Hamiltonian system, it only captures the point spectrum(eigenvalues) and misses the entire spectrum(eigenvalues + continuous spectrum). It also still suffers from spectral pollution and requires careful tuning of the time delay parameter. With 300 basis functions, ResDMD is still unable to fully capture the entire spectrum, i.e., the unit circle, due to the insufficient number of basis functions. In the original ResDMD work, 964 basis functions using a hyperbolic cross approximation of order 100 were required to adequately cover the spectrum with a dataset of the same size [12, Section 4.3.1]. Our method, in contrast, represents the results using shaded areas that show the pseudospectrum, which is a key feature that can capture the whole spectrum. While the shaded area may appear broad, this actually demonstrates our method’s ability to detect the complete spectrum. This radius of the shaded region accounts for computational uncertainties, as exact spectrum computation is computationally impossible. Theoretically, ResDMD guarantees that as this error tolerance(radius of the shaded region) approaches zero, the pseudospectrum converges to the true spectrum (the unit circle in this case) without spectral pollution. This comparison demonstrates that NN-ResDMD, even with only 300 basis functions, outperforms all four classical methods in terms of capturing the complete spectrum with greater accuracy and fewer basis functions.
NN-ResDMD: $N_{K}=25$
NN-ResDMD: $N_{K}=50$
NN-ResDMD: $N_{K}=100$
NN-ResDMD: $N_{K}=300$
Figure 2: The four plots depict the spectrum of the Koopman operator, constructed using varying dictionary size $N_{K}$ of 25, 50, 100, and 300. Each plot utilizes 90 initial points to illustrate the impact of increasing the dictionary size on approximating the spectrum of the Koopman operator.
NN-ResDMD: $N_{K}=50$
NN-ResDMD: $N_{K}=150$
NN-ResDMD: $N_{K}=250$
NN-ResDMD: $N_{K}=350$
Figure 3: Same example as Figure 4 but with larger data size, using 240 initial points to show the effect of increasing dictionary size on approximating the Koopman operator spectrum.
EDMD: $N_{K}=300$
EDMD-DL: $N_{K}=300$
Hankel-DMD: $N_{K}=300$
ResDMD: $N_{K}=300$
Figure 4: Comparison with classical methods. The four plots above represent the spectral information obtained from a $300\times 300$ Koopman matrix, calculated using four methods: EDMD, EDMD with Dictionary Learning (EDMD-DL), Hankel-DMD, and ResDMD. The illustrated eigenvalue spectra of the Koopman operator highlight the differences in results produced by these methods.
4.2 Turbulence
Recovering spatial patterns is a typical goal of DMD-based methods, especially in fluid dynamics, where Kernel ResDMD has been particularly successful in capturing such patterns and detecting acoustic vibrations in the turbulence system. However, Kernel ResDMD requires careful selection of kernel functions, while NN-ResDMD bypasses this by using neural networks to train observables and compute Koopman modes.
We demonstrate this by applying NN-ResDMD to the turbulence system using the dataset from [12, Section 6.3]. The ground truth in the first plot of Figure 5 represents a high-dimensional pressure field distribution (approximately 30,000 spatial dimensions) around an airfoil, with a clear distinction between the upper and lower surfaces. Technically, we apply truncated Singular Value Decomposition (SVD), select 300 observables, compute Koopman modes, and project them back into the original state space.
In Figure 5, the first Koopman mode estimated by NN-ResDMD with the smallest residual value successfully highlights a clear global spatial separation that aligns with the pattern observed in the original pressure field. This advantage allows the first Koopman mode to directly distinguish spatial features present in the true pressure field, making it a powerful tool for interpreting complex fluid dynamics data. Subsequent Koopman modes also reveal strong acoustic waves that are critical in various aeronautical engineering fields. In contrast, Kernel ResDMD with a generic normalized Gaussian kernel function, as shown in the original work, is unable to produce a Koopman mode similar to the first Koopman mode from NN-ResDMD that clearly distinguishes the pressure field. For comparison, we also plot four Koopman modes computed by Hankel-DMD with a time delay of 5, corresponding to the four smallest residual values, which similarly do not reveal the pressure field patterns as in NN-ResDMD. These results are presented in Appendix Figure 7 and Appenxic A.6.2.
Figure 5: The plots illustrate turbulence detection with Koopman modes computed by 300 observables. The first plot shows a 2D scatter plot of the pressure field, while the other plots display various Koopman modes, each labeled with corresponding residuals. The small residual values in the figures associated with the Koopman modes confirm the estimation accuracy.
4.3 Identification of neural dynamics in mice visual cortex
Since NN-ResDMD directly minimizes the residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics. To evaluate how effectively NN-ResDMD reveals latent temporal dynamics in real data, we apply it to a dataset of high-dimensional neural signals and demonstrate its advantages over a series of classical methods: the Hankel-DMD, EDMD (combined with RBF basis) and Kernel ResDMD. These methods are selected as representative approaches for handling high-dimensional data.
The dataset is part of the open dataset on mice from the competition "Sensorium 2023" [26, 27]. In the experiments, mice viewed natural videos while their neural signals were recorded via calcium imaging in the primary visual cortex, reflecting the activity of thousands of neurons. Here, we focus on the state partitioning of neural signals. Specifically, in each mouse, six video stimuli were repeatedly shown, creating ideal conditions to define brain states. Neural activity during repeated trials with the same stimuli is assumed to reflect the same underlying dynamic system, enabling Koopman decomposition methods to uncover and separate these brain states.
The dataset consists of neural recordings from five mice, each exposed to 6 video stimuli, repeated 9-10 times for a total of around 60 trials. Each recording captures the activity of over 7,000 neurons, with each 10-second video sampled at 50 Hz, resulting in 300 data points per trial.
We applied NN-ResDMD and three classical Koopman decomposition methods (Hankel-DMD, EDMD with RBF basis, and Kernel ResDMD) to these datasets, using different implementations and Koopman subspace dimensions. For NN-ResDMD, we trained dictionaries on all snapshots from each mouse to avoid overfitting, reduced the data to 300 dimensions via SVD, and selected 501 eigenfunctions. The decomposed eigenfunctions are shown in Figure 6A(top), with markers indicating ground truth state separations. For Hankel-DMD, we built a Hankel matrix with a delay of 50, producing 50 eigenfunctions per trial. In EDMD with RBF basis, we used the SVD-truncated 300 basis and 1000 RBF functions, resulting in 1301 eigenfunctions. For Kernel ResDMD, we used normalized Gaussians as kernel functions, setting the Koopman subspace dimension to 299 eigenfunctions based on [28]. See Appendix A.7.4 for method details and Appendix A.8 for dictionary size justification. These eigenfunctions, shown in Figure 6A(bottom), Appendix Figure 9A, and Appendix Figure 10A, are compared to the ground truth trial identities.
The Koopman eigenfunctions represent dynamical features corresponding to the video stimuli. To evaluate their effectiveness, we assess how well eigenfunctions of the same stimuli cluster together, distinguishing them from other states. If the eigenfunctions capture key dynamics related to the stimuli, those from trials with the same video should be separable from others. This turns the problem into a clustering task based on the separability of eigenfunctions across different stimuli. Note that averaged trial differences are even visibly clear for the NN-ResDMD case.
Figure 6: NN-ResDMD outperforms Hankel-DMD in identifying latent dynamic structures in neural signals with a dictionary size of 501. (A) (Top) 500 Koopman eigenfunctions estimated by NN-ResDMD across 6 states characterized by different video stimuli in an example mouse. Each trial contains 300 data points (10s at 50Hz). (Bottom) 50 Koopman eigenfunctions approximated by Hankel-DMD, each 50 points long, reflecting the dimension of the Hankel matrix. (B) 2D representation of Koopman eigenfunctions for all tested mice, computed by NN-ResDMD and reduced via Multidimensional Scaling (MDS). Trials of the same state cluster well. (C) Same as (B) but computed with Hankel-DMD, showing no clear state separation. (D) 2D representation of Koopman eigenfunctions for the first mouse, computed by EDMD with an RBF basis. See Appendix Figure 9 for full results. (E) Same as (D) but computed with Kernel ResDMD. See Appendix Figure 10 for full results. (F) Davies-Bouldin Indices (DBIs) evaluating clustering quality across four methods (NN-ResDMD, Hankel-DMD, EDMD+RBF, and Kernel ResDMD) for five mice. Lower DBI values for NN-ResDMD indicate better clustering.
We use Multi-dimensional Scaling (MDS) to visualize how these eigenfunction-based features cluster according to ground truth states. MDS reduces data dimensionality based on similarities, making it ideal for visualizing clustering performance. While UMAP and t-SNE are alternative methods, we show MDS results in 2D space (Figure 6B-E), with similar results for UMAP and t-SNE in the supplementary materials (Appendix Figure 8, Appendix Figure 9C,D and Appendix Figure 10C,D).
The 2D MDS visualization reveals clear separation of features for all 5 mice using NN-ResDMD (Figure 6B), whereas no other method shows clear clustering (Figure 6C-E, Appendix Figure 9B, Appendix Figure 10B). To quantify this clustering, we calculate the Davies-Bouldin index (DBI), a measure of clustering quality that assesses how compact and well-separated the clusters are. A lower DBI indicates more compact clusters that are farther apart from each other, which corresponds to better clustering. The DBI is significantly lower for NN-ResDMD (Figure 6F), suggesting that it captures the latent dynamic structure more effectively than all three other methods. Similar clustering patterns are confirmed with UMAP and t-SNE (Appendix Figure 11).
5 Conclusion and future work
Koopman spectral components (eigenpairs) are fundamental to understanding dynamical systems, as they reveal intrinsic patterns and structures underlying complex temporal behavior through a linear framework for analyzing nonlinear dynamics. In this paper, we introduced NN-ResDMD, a method for estimating eigenpairs by minimizing spectral residuals, eliminating ResDMD’s need to filter precomputed results. Despite higher computational costs, using neural networks to learn eigenpairs provides a significant advantage by capturing patterns automatically and reducing manual intervention in basis selection. This flexibility is particularly beneficial for high-dimensional systems where traditional methods often struggle. Our experiments demonstrate that NN-ResDMD outperforms classical methods—including EDMD, Hankel-DMD, ResDMD, and their variants—in uncovering critical spatiotemporal characteristics of nonlinear dynamics.
Despite the advantages, NN-ResDMD has several limitations and we discuss the major ones here. First, the neural network structure incurs higher computational costs compared to classical approaches, making it unsuitable for real-time learning tasks (see a brief discussion in Appendix A.5). Second, the deterministic nature of the framework does not account for stochastic aspects of the system, such as those addressed by methods like VAMP [29], limiting its applicability to highly noisy data. Additionally, the performance of NN-ResDMD is sensitive to hyperparameter tuning, including network architecture, dictionary size, and training criteria, which can require significant effort to optimize.
Koopman eigenpairs provide unique perspectives into the interpretation of nonlinear dynamical mechanisms, and feedforward neural networks (FNNs) represent an initial step in learning spectral properties directly from data. In recent years, various deep neural network structures have been employed to learn the Koopman representations with different optimization targets other than the spectral residuals (e.g. [30, 21, 29, 31, 22, 32, 33, 34](see Appendix section A.4 for a comparison with the VAMP framework). With our approach, we demonstrate that even basic architectures can achieve significant improvements in Koopman operator estimation by using the spectral residual loss. Therefore, future work could focus on refining neural network architectures to enhance the accuracy and efficiency of Koopman eigenpair estimation. One promising direction is the incorporation of Physics-Informed Neural Networks (PINNs) and Physics-Informed Neural Operators (PINOs), which integrate physical laws directly into the learning process. This integration will ensure that the resulting Koopman eigenfunctions align with known physical constraint, avoid overfitting and faciliatates generalization. Indeed, the integration of PINNs and PINOs with the Koopman framework has the potential to serve as a powerful bridge between data-driven and model-driven approaches, offering enhanced insights into complex systems and enabling more robust temporal evolution predictions.
Acknowledgments
Authors are grateful to Matthew Colbrook for valuable guidance on ResDMD technical part and providing fluid data, to Bin Han for discussion on computational issues, to Michel Beserve and Rory Bufacchi for proofreading and to Turishcheva Polina for providing permission of using the Sensorium 2023 dataset.
Appendix A Appendix
A.1 Calculation steps for 3.3
Here we are going to show how squared relative residual implies (A.1) and then implies (3.3). Consider $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}\mathbf{v}_{i}$ with $\|\phi\|_{2}=1$, then
[…]
$=\langle\mathcal{K}\phi-\lambda\phi,\mathcal{K}\phi-\lambda\phi\rangle_{\mu}$
$=\langle\mathcal{K}\phi,\mathcal{K}\phi\rangle_{\mu}-\langle\lambda\phi,\mathcal{K}\phi\rangle_{\mu}-\langle\mathcal{K}\phi,\lambda\phi\rangle_{\mu}+\langle\lambda\phi,\lambda\phi\rangle_{\mu}$
$=\langle\mathcal{K}\mathbf{\Psi}\mathbf{v},\mathcal{K}\mathbf{\Psi}\mathbf{v}\rangle_{\mu}-\bar{\lambda}\langle\mathbf{\Psi}\mathbf{v},\mathcal{K}\mathbf{\Psi}\mathbf{v}\rangle_{\mu}-\lambda\langle\mathcal{K}\mathbf{\Psi}\mathbf{v},\mathbf{\Psi}\mathbf{v}\rangle_{\mu}+|\lambda|^{2}\langle\mathbf{\Psi}\mathbf{v},\mathbf{\Psi}\mathbf{v}\rangle_{\mu}$
$=\langle\sum_{i=1}^{N_{K}}\mathcal{K}\psi_{i}\mathbf{v}_{i},\sum_{j=1}^{N_{K}}\mathcal{K}\psi_{j}\mathbf{v}_{j}\rangle_{\mu}-\bar{\lambda}\langle\sum_{i=1}^{N_{K}}\psi_{i}\mathbf{v}_{i},\sum_{j=1}^{N_{K}}\mathcal{K}\psi_{j}\mathbf{v}_{j}\rangle_{\mu}-\lambda\langle\sum_{i=1}^{N_{K}}\mathcal{K}\psi_{i}\mathbf{v}_{i},\sum_{j=1}^{N_{K}}\psi_{j}\mathbf{v}_{j}\rangle_{\mu}+|\lambda|^{2}\langle\sum_{i=1}^{N_{K}}\psi_{i}\mathbf{v}_{i},\sum_{j=1}^{N_{K}}\psi_{j}\mathbf{v}_{j}\rangle_{\mu}$
$=\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}\mathbf{v}_{j}-\bar{\lambda}\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}\mathbf{v}_{j}-\lambda\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}\mathbf{v}_{j}+|\lambda|^{2}\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\langle\psi_{i},\psi_{j}\rangle_{\mu}\mathbf{v}_{j}$
[…]
$\approx\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\left[\frac{1}{m}[\Psi_{Y}^{*}\Psi_{Y}]_{ij}-\bar{\lambda}\frac{1}{m}[\Psi_{X}^{*}\Psi_{Y}]_{ij}-\lambda\frac{1}{m}[\Psi_{Y}^{*}\Psi_{X}]_{ij}+|\lambda|^{2}\frac{1}{m}[\Psi_{X}^{*}\Psi_{X}]_{ij}\right]\mathbf{v}_{j}$
[…]
Remark.
[…]
A.2 Details for deriving (3.5)
$J$ $=\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}$
$=\sum_{i=1}^{N_{K}}\frac{1}{m}\mathbf{v}_{i}^{*}\left[\Psi_{Y}^{*}\Psi_{Y}-\lambda_{i}(\Psi_{X}^{*}\Psi_{Y})^{*}-\bar{\lambda}_{i}\Psi_{X}^{*}\Psi_{Y}+|\lambda_{i}|^{2}\Psi_{X}^{*}\Psi_{X}\right]\mathbf{v}_{i}$
$=\sum_{i=1}^{N_{K}}\frac{1}{m}\left[\mathbf{v}_{i}^{*}(\Psi_{Y}^{*}\Psi_{Y})\mathbf{v}_{i}-\mathbf{v}_{i}^{*}(\Psi_{X}^{*}\Psi_{Y})^{*}\lambda_{i}\mathbf{v}_{i}-\mathbf{v}_{i}^{*}\bar{\lambda}_{i}(\Psi_{X}^{*}\Psi_{Y})\mathbf{v}_{i}+\mathbf{v}_{i}^{*}K^{*}(\Psi_{X}^{*}\Psi_{X})K\mathbf{v}_{i}\right]$
$=\sum_{i=1}^{N_{K}}\frac{1}{m}\left[\mathbf{v}_{i}^{*}(\Psi_{Y}^{*}\Psi_{Y})\mathbf{v}_{i}-\mathbf{v}_{i}^{*}(\Psi_{X}^{*}\Psi_{Y})^{*}K\mathbf{v}_{i}-\mathbf{v}_{i}^{*}K^{*}(\Psi_{X}^{*}\Psi_{Y})\mathbf{v}_{i}+\mathbf{v}_{i}^{*}K^{*}(\Psi_{X}^{*}\Psi_{X})K\mathbf{v}_{i}\right]$
$=\sum_{i=1}^{N_{K}}\frac{1}{m}\bigg{(}\langle\Psi_{Y}\mathbf{v}_{i},\Psi_{Y}\mathbf{v}_{i}\rangle-\langle\Psi_{Y}\mathbf{v}_{i},\Psi_{X}K\mathbf{v}_{i}\rangle$
$\qquad-\langle\Psi_{X}K\mathbf{v}_{i},\Psi_{Y}\mathbf{v}_{i}\rangle+\langle\Psi_{X}K\mathbf{v}_{i},\Psi_{X}K\mathbf{v}_{i}\rangle\bigg{)}$
$=\sum_{i=1}^{N_{K}}\frac{1}{m}\langle\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i},\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i}\rangle$
$=\sum_{i=1}^{N_{K}}\frac{1}{m}\|\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i}\|_{2}^{2}$
$=\frac{1}{m}\|(\Psi_{Y}-\Psi_{X}K)V\|_{F}^{2}.$
Next, by matrix calculus with denominator layout convention, we try to find minimal of $J$:
$0=\frac{dJ}{dK}$ $=\frac{d\operatorname{tr}(J)}{dK}\quad(\text{since $J$ is a scalar})$
$=\frac{d}{dK}\operatorname{tr}\bigg{(}\frac{1}{m}\sum_{i=1}^{N_{K}}\mathbf{v}_{i}^{*}\bigg{[}\Psi_{Y}^{*}\Psi_{Y}-(\Psi_{X}^{*}\Psi_{Y})^{*}K$
$\qquad-K^{*}(\Psi_{X}^{*}\Psi_{Y})+K^{*}(\Psi_{X}^{*}\Psi_{X})K\bigg{]}\mathbf{v}_{i}\bigg{)}$
$=\sum_{i=1}^{N_{K}}\frac{d}{dK}\operatorname{tr}\bigg{(}\mathbf{v}_{i}^{*}\bigg{[}L-A^{*}K-K^{*}A+K^{*}GK\bigg{]}\mathbf{v}_{i}\bigg{)}$
$=\sum_{i=1}^{N_{K}}\frac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}L\mathbf{v}_{i}\right)+\frac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}A^{*}K\mathbf{v}_{i}\right)+\frac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}A\mathbf{v}_{i}\right)+\frac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)$
$=\sum_{i=1}^{N_{K}}-A\mathbf{v}_{i}\mathbf{v}_{i}^{*}-A\mathbf{v}_{i}\mathbf{v}_{i}^{*}+(G+G^{*})K\mathbf{v}_{i}\mathbf{v}_{i}^{*}$
$=\sum_{i=1}^{N_{K}}(-2A+2GK)\mathbf{v}_{i}\mathbf{v}_{i}^{*}\quad(\text{$G$ is symmetric})$
where $\operatorname{tr}()$ is trace of a matrix and $G=\Psi_{X}^{*}\Psi_{X},A=\Psi_{X}^{*}\Psi_{Y},L=\Psi_{Y}^{*}\Psi_{Y}.$
Since eigenvector $v_{i}$ is not a zero vector, $v_{i}v_{i}^{*}$ is not a zero matrix. So
$-2A+2GK=0\Rightarrow K=G^{\dagger}A.$
Remark.
To solve $\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)$ , we simply rewrite it as
$\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)=\tfrac{d}{dK}\operatorname{tr}\left((K\mathbf{v}_{i})^{*}G(K\mathbf{v}_{i})\right)$
A.3 Discussion On Convergence
To understand how neural networks enhance NN-ResDMD, it is important to introduce Barron space [35, 36, 23, 37]. Barron space characterizes functions efficiently approximated by two-layer neural networks, which is central to NN-ResDMD. By leveraging networks that approximate functions within this space, NN-ResDMD can flexibly optimize the dictionary functions for Koopman operator approximation, making it highly effective for complex, high-dimensional systems.
A function $f$ belongs to Barron space $\mathcal{B}$ if it can be represented as:
$f(x)=\int_{\Omega}a\sigma(w^{T}x)\rho(da,dw),$
where $\sigma$ is the activation function, $w$ is a weight vector, $a$ is a coefficient, and $\rho$ is a probability distribution. The complexity of $f$ is measured by the Barron norm $\|f\|_{\mathcal{B}}$:
$\|f\|_{\mathcal{B}}=\inf_{\rho\in P_{f}}\left(\int_{\Omega}|a|\|w\|_{1}\rho(da,dw)\right),$
where $P_{f}$ is the set of distributions for which $f$ can be represented. This framework provides a basis for analyzing approximation errors in neural networks.
The following theorem [38] discusses the approximation capabilities of two-layer neural networks within this context, establishing a foundation for the subsequent analysis.
Theorem A.1 (Direct Approximation Theorem, $L^{2}$-version).
For any $f\in\mathcal{B}$ and $r\in\mathbb{N}$, there exists a two-layer neural network $f_{r}$ with $r$ neurons $\{(a_{i},\mathbf{w}_{i})\}$ such that
$\|f-f_{r}\|_{L^{2}}\lesssim\frac{\|f\|_{\mathcal{B}}}{\sqrt{r}}.$
This result implies that the approximation error decreases at a rate of $1/\sqrt{r}$ as the number of neurons $r$ increases, with the constant $\|f\|_{\mathcal{B}}$ reflecting the complexity of the function $f$ within the Barron space.
Now, consider a Barron space $\mathcal{B}$ which is dense in $L^{2}(\Omega,\mu)$ and a projected Koopman operator $\mathcal{K}_{N_{K}}:\mathcal{B}_{N_{K}}\to L^{2}(\Omega,\mu)$ where $\mathcal{B}_{N_{K}}\subseteq\mathcal{B}$ is a $N_{K}$-dimensional subspace spanned by some dictionary $\mathbf{\Psi}=\{\psi_{i}\}_{i=1}^{N_{K}}$. According to Theorem A.1, we can have a well-trained dictionary that almost spans $\mathcal{B}_{N_{K}}$, i.e., given $\epsilon>0$, we can always obtain a dictionary $\mathbf{\Psi}_{r}=\{\psi_{r,i}\}_{i=1}^{N_{K}}$ such that $\sum_{i=1}^{N_{K}}\|\psi_{r,i}-\psi_{i}\|^{2}_{2}<\epsilon$.
A.4 Highlights of NN-ResDMD compared with typical existing neural network-based Koopman framework
Our NN-ResDMD method takes a fundamentally different approach from existing deep learning methods by building upon the residual-based framework of ResDMD rather than the different Koopman-approximating loss functions following the variational principles of VAMPnets [39, 40, 29] or the deep autoencoder structure in [41]. By incorporating spectral residual measures into deep learning and introducing a structured representation that captures dependencies among eigenvalues, we achieve more compact and interpretable models for nonlinear systems with continuous spectra. This approach enables us to directly minimize Koopman spectral approximation errors while avoiding the high-dimensional representations or point-spectrum limitations of previous methods.
If we take the VAMP framework as an example, here are the connections and differences. The proposed loss function and the VAMP score share the goal of optimizing approximations of the Koopman operator’s spectral properties, establishing a connection in their ultimate purpose. However, although they both depend on the covariance matrices (in our manuscript Equation 3.2), their methodologies differ significantly. Our residual-based method directly minimizes the spectral approximation error of the Koopman operator and accommodates both point and continuous spectra, while the VAMP score follows a variational framework, maximizing the sum of singular values to approximate the point spectrum, primarily for stochastic systems (though see an exception in TianWu+2021+635+659 [39]). Moreover, while VAMP is specifically designed for Markov processes and requires the Koopman operator to be Hilbert-Schmidt, our approach focuses on deterministic systems and enables a more comprehensive spectral analysis that incorporates continuous spectra. This distinction in scope and methodology highlights how the two frameworks complement each other in addressing different aspects of spectral estimation.
A.5 Discussion of computation costs
Despite the various advantages of the NN-ResDMD framework, one significant limitation is its higher computational cost compared to the original ResDMD and other classical methods.
Theoretical Perspective: The NN-ResDMD algorithm’s computational demands stem primarily from its iterative optimization process. Each iteration involves a gradient descent update with complexity scaling linearly with both system dimensionality and neural network parameters. Though individual gradient steps are computationally lightweight for standard network architectures, the algorithm’s efficiency issue lies in its repeated least-squares optimizations. Below we provide a detailed comparison between computation costs of several methods we used in this paper.
Comparison over methods: The computational costs of EDMD, ResDMD, EDMD-DL, Hankel-DMD, and NN-ResDMD vary significantly based on their core computational steps and specific configurations. For a dataset with $m=10^{5}$ data points and $N_{K}=300$ dictionary functions (for EDMD-based methods), the theoretical complexity and runtime differ across methods. EDMD involves least squares and eigenvalue decomposition, with a complexity of $O(N_{K}^{2}m+N_{K}^{3})$, making it the fastest method, and typically requiring only seconds to minutes for computation. ResDMD extends EDMD by adding residual evaluation and pseudospectrum computation. The residual evaluation introduces an additional $O(N_{K}^{3})$, and pseudospectrum computation across $n_{z}$ grid points incurs $O(n_{z}N_{K}^{3})$, resulting in a total complexity of $O(N_{K}^{2}m+n_{z}N_{K}^{3})$. This leads to runtimes ranging from minutes to hours, depending on the resolution of the pseudospectrum grid. EDMD-DL incorporates dictionary learning through stochastic gradient descent (SGD), where each iteration involves matrix construction ($O(N_{K}^{2}m)$), Koopman matrix computation ($O(N_{K}^{3})$), and neural network forward/backward propagation ($O(d|H|)$, with $d|H|$ representing the total network parameter size). With $k$ SGD iterations, the total complexity becomes $O(k(N_{K}^{2}m+N_{K}^{3}+d|H|))$, leading to runtimes also in the range of minutes to hours depending on $k$. NN-ResDMD, which builds on EDMD-DL, shares the same complexity, $O(k(N_{K}^{2}m+N_{K}^{3}+d|H|))$, but includes the explicit use of Koopman matrix eigenvectors and optional pseudospectrum computation, making its runtime slightly longer than EDMD-DL for high-resolution spectral analysis. Hankel-DMD, using a time delay embedding dimension $T$, constructs a Hankel matrix ($O(Tm)$), performs singular value decomposition (SVD) ($O(T^{2}m)$), and computes the eigenvalues of a reduced $T\times T$ matrix ($O(T^{3})$). The total complexity is $O(Tm+T^{2}m+T^{3})$, and the runtime is heavily influenced by $T$, typically ranging from minutes to hours. While EDMD is computationally the most efficient, ResDMD and Hankel-DMD provide higher precision and robustness in spectral analysis at the expense of increased runtime, and EDMD-DL and NN-ResDMD offer flexibility and accuracy through dictionary learning, with additional SGD iterations and optional pseudospectrum computation contributing to their computational burden.
Empirical Perspective: In our experiments, without computing the pseudospectrum, the computational cost of ResDMD typically ranges from seconds to minutes. NN-ResDMD, on the other hand, can require tens of minutes to several hours, depending on factors such as data dimensionality, the number of snapshots, hidden layer configurations, dictionary sizes, and training convergence criteria.
Trade-off Between Cost and Accuracy: While NN-ResDMD’s additional computational steps introduce higher costs, they enhance the accuracy and robustness of Koopman eigenpair estimation by allowing automatic dictionary learning and minimizing spurious spectral components. This trade-off makes NN-ResDMD particularly valuable in applications where precision is critical. However, its computational demands render it less suitable for real-time or online Koopman model learning tasks.
A.6 Hankel-DMD
A.6.1 Justification of Using Hankel-DMD as comparison in all experiments
Hankel-DMD operates by constructing a Hankel matrix from time-delayed measurements of the system state, based on Takens’ embedding theorem, which states that time-delayed coordinates can reconstruct the state space of dynamical systems. Hankel-DMD also falls within the framework of Extended Dynamic Mode Decomposition (EDMD), as it effectively uses time-delayed states as dictionary functions. This connection introduces convergence conditions specific to time-delay embeddings, differing from those associated with standard EDMD implementations. This makes Hankel-DMD a natural choice for comparison in the pendulum system. Specifically, the method enables a more detailed extraction of the system’s modes and dynamics, with theoretical guarantees established in works like [25], which proved its convergence for ergodic systems.
Practically, the approach involves constructing a large matrix of time-shifted copies of measured data, where the number of delays determines how many past states are considered. This theoretically grounded framework is particularly effective when the system states have good temporal resolution and has shown strong performance in analyzing high-dimensional dynamical systems. Consequently, we also apply Hankel-DMD to the turbulence and neural dynamics experiments to evaluate its effectiveness in these representative high-dimensional settings.
As results, although its performance rivals NN-ResDMD in the simple pendulum system by showing eigenvalues points near the unit circle and containing some polluted eigenvalues, which are close to the ground truth unit circle, we would like to emphasize that it capture the point spectrum and miss the full spectral information. When it comes to high-dimensional systems, it fails to capture key dynamics in higher-dimensional systems, as seen in the later experiment (Section 4.2 and Section 4.3).
A.6.2 Application in Turbulence
Here we present the Koopman modes computed by Hankel-DMD for comparison with the NN-ResDMD results. As shown in Figure 7, despite having small residuals, these modes fail to clearly capture the fundamental pressure field structure that was successfully identified by NN-ResDMD’s first Koopman mode (see Figure 5). This comparison demonstrates the superior ability of NN-ResDMD to extract physically meaningful patterns from complex fluid systems.
Figure 7: The plots illustrate turbulence detection using the four Koopman modes computed by Hankel-DMD, which are ranked with their corresponding residuals from the smallest.
A.7 Practical details for neural data analysis
A.7.1 Dataset Details and Experimental Setup
The dataset utilized in this study is part of the open dataset provided for the ’Sensorium 2023’ competition [26]. The dataset consists of calcium imaging recordings from the primary visual cortex of mice. During the experiments, the mice were presented with natural video stimuli while the activity of thousands of neurons was recorded. The objective of the competition is to predict large-scale neuronal population activity in response to different frames of the stimulus videos, based on the hypothesis that population dynamics in the primary visual cortex, driven by visual stimuli, encode significant information about the dynamics of the videos [42, 43, 44].
A.7.2 Task Definition and Rationale
In contrast to the competition’s prediction objective, our study focuses on the task of state partitioning of neural signals. While prediction remains feasible, we aim to demonstrate that state partitioning is sufficient to highlight the superiority of NN-ResDMD over a series of other methods in uncovering the latent dynamics of the system. Specifically, in each experiment, a set of six video stimuli was repeatedly presented to each mouse, creating ideal conditions for defining brain states. The recording setup remained consistent for each mouse, ensuring that the neural activities could be interpreted as originating from the same dynamical system, with the primary variable being the input stimulus.
We hypothesize that during repeated trials with identical visual stimuli, the underlying dynamics of the neural system remain consistent. Consequently, the recurrence of the same brain state is expected during these trials. This provides a reliable basis for testing the efficacy of Koopman decomposition methods in uncovering latent dynamics and distinguishing these states.
A.7.3 Dataset Structure and Dimensionality
The dataset includes neural recordings from five mice, with each mouse responding to six distinct video stimuli, presented in 9-10 repeated trials (resulting in approximately 60 trials in total). Each trial involves recordings of over 7000 neurons. The duration of each video stimulus is 10 seconds, with a sampling rate of 50 Hz, yielding 300 data points (299 snapshots) per trial. Thus, the data to be analyzed consists of a high-dimensional time series with 7000+ observables per snapshot.
A.7.4 Implementations of NN-ResDMD and other classical methods
We compare here four methods: the proposed NN-ResDMD and three classical Koopman decomposition methods for high-dimensional systems: the Hankel-DMD, the EDMD with RBF basis, and the Kernel ResDMD. We applied them to the 5 datasets, although with slightly different implementations and different dimensions of approximated Koopman invariance subspace.
For NN-ResDMD, we train the dictionaries with all the snapshots recorded in each mouse such that the total snapshot number is the product of the snapshot number in one trial and the number of all trials. This is to avoid overfitting with the small snapshot numbers within a trial. The high-dimensional data is first reduced to 300 dimensions with Singular Value Decomposition. The dimension of the Koopman subspace is chosen to be 601, consisting of 300 trained bases and 301 pre-chosen ones (constant and the first-degree polynomials of the SVD-ed 300 dimensions). The first 501 eigenfunctions sorted by the modulus of eigenvalues are selected to avoid spurious eigenvalues estimation due to noise. One can find the decomposed eigenfunctions in Figure 6A(top), with a marker of the ground truth state separations based on stimulus identity.
For Hankel-DMD, the Koopman eigenfunctions were approximated using the eigenvectors of the Hankel matrix. Specifically, the Hankel matrix was formed as in Equation 53 from arbabi2017ergodic [25], using all the observables from one trial of each mouse with a delay of 50. Consequently, the snapshot size became 249 times the observable number, and the resulting number of eigenfunctions was 50, each with a length of 50. The Hankel-DMD eigenfunctions for each trial of data are shown in Figure 6A (bottom), alongside the ground truth trial identities for comparison.
For EDMD with RBF basis, the high-dimensional dataset is first reduced to 300 dimensions with SVD. Then RBF basis is calculated with 1000 RBF functions. The choice of the basis number is decided based on classical experiments of using RBF basis to estimate the Koopman operator of Duffing systems [18].
For Kernel ResDMD, as it is a variant of Kernel EDMD [45], the dimension of the Koopman invariant subspace should corresponds to the sample number (in time). Given the data size to be 300, we have 299 snapshots, resulting in 299 Koopman bases. The detailed calculated is performed for each trial with the program provided in the original ResDMD paper [28, 12]. We chose the kernel function as the commonly-used normalized Gaussian function in the calculation.
The Koopman eigenfunctions from both NN-ResDMD and other methods represent dynamical features corresponding to one of the six video stimuli. To evaluate how well the eigenfunctions capture the latent dynamics, we assess the similarity of the features for trials with the same stimulus and their dissimilarity from those corresponding to different stimuli. Effectively, this makes the problem a clustering task, where the separability of the Koopman eigenfunctions reflects how well they capture the key dynamic components related to the stimuli.
A.8 Choice justification of dictionary sizes
In this section, we provide justifications for the use of different dictionary sizes (i.e., the number of Koopman eigenfunctions) in the aforementioned four methods for the neural dynamics experiment.
First, the high-dimensional data was pre-processed using SVD to reduce its dimensionality to 300. Then, for the four methods:
  1. 1.
    For NN-ResDMD, we selected 300 trained basis functions and 300 first-order monomial basis functions as the dictionary for the 300 reduced observables. This choice ensures the dictionary is rich enough to span the Koopman invariant subspace. Hence, the size of the trained dictionary was set to be at least equal to the original observable size. Then based on the rank of estimated Koopman eigenvalues, we select the dominant 501 eigenfunctions to avoid the eigenfunctions with zero eigenvalues.
  2. 2.
    For Hankel DMD, the number of delays (as dictionary size/number of eigenfunctions) is first constrained by the temporal sample size (i.e., snapshot size) because it cannot exceed the maximum snapshot size. Therefore, it is impossible to choose the same dictionary size as the NN-ResDMD example. Choosing the delay too small will result in an insufficient dictionary size to span the Koopman invariant subspace, and too large will reduce the actual snapshot size to estimate the covariance matrices in the estimation of the Koopman matrix. Therefore, we chose a compromise delay number of 50 that satisfies both needs.
  3. 3.
    For RBF basis, in principle, we can use the same dictionary size. However, our previous experience with a similar dataset and the results of using the RBF basis for the EDMD method all suggest that the performance will be better with more dictionary functions. Therefore, we chose 1000 RBF basis and the original 300 first-order monomial basis as a better condition compared to the same dictionary size with NN-ResDMD.
  4. 4.
    For Kernel ResDMD, the dictionary size is theoretically determined to be the number of snapshots. Therefore, we cannot make the dictionary size consistent with the NN-ResDMD example.
Based on the above justifications, we believe our choices of dictionary sizes are reasonable and ensure a fair comparison across the methods.
A.8.1 Visualization and Clustering Performance
To visualize the clustering of high-dimensional Koopman eigenfunctions, we perform dimensionality reduction using Multi-dimensional Scaling (MDS). MDS is particularly useful for visualizing high-dimensional data by preserving pairwise similarities [46] (here we use correlation as a measure of similarities). While UMAP [47] and t-SNE [48] are alternative visualization methods, with different emphasis on global-local relationships, we primarily use MDS in this study and provide UMAP and t-SNE results in the supplementary materials (see Appendix Figure 8A, B, Appendix Figure 9C, D and Appendix Figure 10C, D). UMAP in implementation is still correlation-based. For t-SNE estimation we use the perplexity of 15, as a value for optimal separation.
By applying MDS, the high-dimensional eigenfunction-based features are reduced to a low-dimensional space. For illustration, we present the results of reducing the feature space to two dimensions (Figure 6B-E). The NN-ResDMD reduced features for the six types of trials (corresponding to the six video stimuli) are well-separated for all five mice (Figure 6B). In contrast, the Hankel-DMD features show no clear clustering structure (Figure 6C). Similarly, the features produced by EDMD with an RBF basis and Kernel ResDMD do not show clear separability (Figure 6D-E, Appendix Figure 9B-D, Appendix Figure 10B-D).
Figure 8: State Partition performance of eigenfunctions for NN-ResDMD and Hankel-DMD in 2D space visualized with UMAP (A) and t-SNE (B).
Figure 9: Full results of EDMD with RBF basis. (A) 1301 Koopman eigenfunctions estimated by EDMD with RBF basis in 6 states characterized by 6 different video stimuli in an example mouse. Eigenfunctions in each trial of each state contain 300 data points (10s with a sampling rate of 50Hz). (B) 2-D representation of Koopman eigenfunctions for each trial of all tested mice, calculated by EDMD with RBF basis and reduced by Multidimensional Scaling (MDS). No clear separation of states can be seen from the reduced representation. (C) Same as (B) but visualized with UMAP. No clear separation of states can be seen from the reduced representation. (D) Same as (C) but visualized with t-SNE. No clear separation of states can be seen from the reduced representation.
Figure 10: Same as Figure 9 but estimated with Kernel ResDMD, with 299 basis of the Koopman subspace, thus 299 eigenfunctions.
Figure 11: Davies-Bouldin Indices evaluating the clustering performance of dynamical components learned by four methods (NN-ResDMD, Hankel DMD, EDMD+RBF, and Kernel ResDMD) across five mice. Comparisons are shown using UMAP (A) and t-SNE (B).
A.8.2 Clustering Quality Metrics
We further quantified the clustering quality by calculating the Davies-Bouldin Index (DBI) for both Koopman decomposition methods across all mice (Figure 6F). The DBI is designed to assess the compactness of clusters and the separability between them. A lower DBI indicates better clustering performance. NN-ResDMD features yield significantly lower DBI scores compared to other methods, confirming that NN-ResDMD produces more clearly defined clusters corresponding to the ground truth trials. Similar clustering results are observed with UMAP and t-SNE (see Appendix Figure 11), further supporting the superior performance of NN-ResDMD in capturing the latent dynamic structure compared to the other classical methods.
A.9 Source code
For reproducibility, the source code will be available at the following anonymous URL: https://anonymous.4open.science/r/ICLR-7305-PROJ. A full version of the codebase will be released upon acceptance of the paper.

### v3

@@
-NN-ResDMD: Learning Koopman Representations for Complex Dynamics with Spectral Residuals
+ResKoopNet: Learning Koopman Representations for Complex Dynamics with Spectral Residuals
 Abstract
-Analyzing long-term behaviors in high-dimensional nonlinear dynamical systems remains a significant challenge.
-The Koopman operator framework has emerged as a powerful tool to address this issue by providing a globally linear perspective on nonlinear dynamics.
-However, existing methods for approximating the Koopman operator and its spectral components, particularly in large-scale systems, often lack robust theoretical guarantees.
-Residual Dynamic Mode Decomposition (ResDMD) introduces a spectral residual measure to assess the convergence of the estimated Koopman spectrum, which helps filter out spurious spectral components.
-Nevertheless, it depends on pre-computed spectra, thereby inheriting their inaccuracies.
-To overcome its limitations, we introduce the Neural Network-ResDMD (NN-ResDMD), a method that directly estimates Koopman spectral components by minimizing the spectral residual.
-By leveraging neural networks, NN-ResDMD automatically identifies the optimal basis functions of the Koopman invariant subspace, eliminating the need for manual selection and improving the reliability of the analysis.
-Experiments on physical and biological systems demonstrate that NN-ResDMD significantly improves both accuracy and scalability, making it an effective tool for analyzing complex dynamical systems.
+Analyzing long-term behaviors in high-dimensional nonlinear dynamical systems remains challenging, with the Koopman operator framework providing a powerful global linearization approach, though existing methods for approximating its spectral components often suffer from theoretical limitations and reliance on predefined dictionaries.
+While Residual Dynamic Mode Decomposition (ResDMD) introduced the spectral residual to assess the accuracy of Koopman operator approximation, it only filters precomputed spectra, which prevents it from fully discovering the Koopman operator’s complete spectral information (a limitation sometimes referred to as the ‘spectral inclusion’ problem).
+We introduce ResKoopNet (Residual-based Koopman-learning Network), a novel method that addresses this limitation by explicitly minimizing the spectral residual to compute Koopman eigenpairs, which can identify a more precise and complete spectrum of the Koopman operator.
+This approach provides theoretical guarantees while maintaining computational adaptability through a neural network implementation.
+Experiments on physical and biological systems demonstrate ResKoopNet’s superior accuracy in spectral approximation compared to existing methods, particularly for systems with continuous spectra and high dimensional, which makes it as an effective tool for analyzing complex dynamical systems.
 1 Introduction
-In the study of complex dynamical systems, a critical challenge lies in accurately extracting and analyzing long-term behavior in high-dimensional nonlinear systems.
-Various data-driven methods [1, 2, 3, 4, 5, 6] have been developed to address this challenge, with the Koopman operator [7, 8] framework emerging as a powerful tool due to its ability to globally linearize nonlinear systems.
-Unlike local linearization methods [9, 10], which approximate dynamics near fixed points, the Koopman operator transforms the entire system into a linear form within an infinite-dimensional space, which allows the use of spectral analysis techniques to study complex dynamics.
+In the analysis of complex dynamical systems, a fundamental challenge is accurately capturing and understanding the long-term behavior of highly nonlinear systems, particularly those with high dimensionality.
+Various data-driven methods [6, 32, 43, 34, 38, 22, 27, 46, 16] have been developed to address this challenge.
+Among these approaches, the Koopman operator framework [18, 19] has emerged as a powerful tool due to its ability to globally linearize nonlinear systems.
+Unlike local linearization methods [13, 12], which only approximate dynamics near fixed points, the Koopman operator transforms the entire system into a linear form within an infinite-dimensional function space, enabling the application of spectral analysis techniques to study complex dynamics.
 Despite its promise, practical computational challenges arise from the infinite-dimensional nature of the Koopman operator.
-Numerical methods such as Extended Dynamic Mode Decomposition (EDMD) [11] have been developed to approximate the Koopman operator using a finite set of observables, making it possible to extract dynamic modes from data.
-However, EDMD lacks theoretical guarantees of convergence and may fail to capture the full Koopman spectrum accurately, particularly in large-scale, complex systems.
-To address these limitations, Residual Dynamic Mode Decomposition (ResDMD) [12] was introduced, offering convergence guarantees through a spectral residual measure that quantifies how well the estimated Koopman spectrum converges to the true spectrum.
-By assessing convergence, ResDMD eliminates spurious spectral components that do not represent the system’s true dynamics, enhancing the reliability of spectral estimation.
-However, ResDMD primarily filters precomputed spectra rather than directly approximating Koopman spectra, limiting its ability to independently refine spectral estimates.
-In this paper, we propose Neural Network-ResDMD (NN-ResDMD), which overcomes this limitation by providing a method to directly compute Koopman eigenpairs by minimizing the spectral residual.
-Additionally, NN-ResDMD employs neural networks to automatically select basis functions, eliminating the need for manual intervention, a common challenge in EDMD-based methods.
-Through experiments on both toy models and real-world high-dimensional systems, we demonstrate that NN-ResDMD significantly improves accuracy and scalability, making it a practical and effective tool for analyzing complex dynamical systems.
+Numerical methods such as Extended Dynamic Mode Decomposition (EDMD) have been developed to approximate the Koopman operator using a finite set of observables, which makes it possible to extract dynamic modes from data.
+Important convergence results for EDMD were established by Korda and Mezić [20], which proved that as the number of samples increases, EDMD converges to the $L^{2}$-orthogonal projection of the Koopman operator on a finite-dimensional subspace.
+They further showed convergence in the strong operator topology as the dimension of the subspace increases, and that accumulation points of the spectra of EDMD correspond to eigenvalues of the Koopman operator with associated eigenfunctions converging weakly, provided the weak limit is nonzero.
+However, despite these theoretical advances, EDMD faces notable limitation in spectral approximation.
+It suffers from spectral pollution, where discretization introduces spurious eigenvalues unrelated to the true operator.
+Moreover, EDMD struggles to accurately capture continuous spectra, which are crucial for characterizing chaotic and complex systems.
+As a result, EDMD can miss critical information about the underlying dynamics, particularly in systems exhibiting rich, complex behavior.
+To address these limitation, Residual Dynamic Mode Decomposition (ResDMD) [8] was introduced to offer convergence guarantees through a spectral residual that measures how well the estimated Koopman spectra are.
+By assessing convergence of spectral residual, ResDMD eliminates spurious spectral components that do not represent the system’s true dynamics, which enhances the reliability of spectral estimation.
+ResDMD also introduces the concept of pseudospectrum, regions where the spectral content is significant even without discrete eigenvalues, which is particularly valuable for analyzing systems with continuous spectra.
+However, ResDMD only filters precomputed polluted spectra, which limits its ability to independently refine spectral estimates.
+In this paper, we propose Residual-based Koopman-learning Network (ResKoopNet), which overcomes these limitations by minimizing spectral residuals over eigenpairs while exploring the optimal dictionary space.
+ResKoopNet employs neural networks to optimize dictionary functions for the Koopman invariant subspace, which enables both discrete eigenvalue computation and pseudospectrum analysis.
+Through experiments on both simulation models and high-dimensional real-world systems, we demonstrate its superior accuracy and scalability over other methods.
 2 Preliminary on Koopman Operator
@@
 $x_{k+1}=F(x_{k}),\quad k\in\mathbb{Z}^{+}.$
-The Koopman operator $\mathcal{K}$ acts on observables $g\in L^{2}(\Omega,\mu)$ as:
-$\mathcal{K}g=g\circ F.$
-Although $F$ is nonlinear, the Koopman operator $\mathcal{K}$ is linear, enabling spectral analysis of the system in the infinite-dimension function space.
-A key aspect of modern Koopman operator theory is Koopman Mode Decomposition (KMD) [6], which represents system dynamics through its spectral components, i.e.
-the eigenvalues, Koopman modes, and eigenfunctions.
-The discrete spectrum is particularly important for insights into long-term behavior, such as periodicity and stability.
-Our analysis emphasizes these spectral components derived from KMD.
-Specifically, we seek eigenpairs $(\lambda_{i},\phi_{i})$, where $\lambda_{i}$ are eigenvalues and $\phi_{i}$ are the corresponding Koopman eigenfunctions.
-One of the most prominent numerical methods to approximate the Koopman operator and its spectral components is the Extended Dynamic Mode Decomposition (EDMD) method, introduced by [11].
-In EDMD, a set of observables (dictionary or basis functions) $\mathbf{\Psi}=[\psi_{1},\dots,\psi_{N_{K}}]$ is selected, and the span of these observables defines the subspace $V_{N_{K}}\coloneqq\text{span}\{\psi_{i}\}_{i=1}^{N_{K}}$.
-Snapshots of the system’s state are then collected, and the method constructs a finite-dimensional approximation of the Koopman operator by solving a least-squares problem that relates the snapshots of observables.
-This enables the computation of eigenvalues, eigenfunctions, and Koopman modes.
-Note that while common choices of dictionary functions are polynomials, Fourier basis, RBF functions, etc., the optimal choice of basis functions is usually unknown a priori and depends heavily on the specific dynamical system.
-Given independent and identically distributed data snapshots $\{(x_{i},y_{i})\}_{i=1}^{m}$ with $y_{i}=F(x_{i})$, two matrices $\Psi_{X}$ and $\Psi_{Y}$ are formed by evaluating the dictionary on the data snapshots:
-$\Psi_{X}=\begin{bmatrix}\psi_{1}(x_{1})&\dots&\psi_{N_{K}}(x_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(x_{m})&\dots&\psi_{N_{K}}(x_{m})\end{bmatrix},\quad\Psi_{Y}=\begin{bmatrix}\psi_{1}(y_{1})&\dots&\psi_{N_{K}}(y_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(y_{m})&\dots&\psi_{N_{K}}(y_{m})\end{bmatrix}.$
-EDMD computes the Koopman matrix approximation as $K=\Psi_{X}^{\dagger}\Psi_{Y}$, where $\Psi_{X}^{\dagger}$ is the pseudo-inverse of $\Psi_{X}$.
-The eigenvalues of $K$ provide approximations of the Koopman operator’s spectrum, and the Koopman eigenfunctions $\phi_{i}$ are approximated as $\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i}$, where $\mathbf{v}_{i}\in\mathbb{C}^{N_{K}}$ is the $i$-th eigenvector of $K$.
+The Koopman operator $\mathcal{K}$ acts on observable $g\in\mathcal{F}$ as:
+$\mathcal{K}g=g\circ F,$
+[…]
+$\quad\quad g(x_{n})=[\mathcal{K}^{n}g](x_{0})$
+[…]
+where $\varphi_{\lambda}$ are the eigenfunctions of $\mathcal{K}$, $c_{\lambda}$ are expansion coefficients, and $\phi_{\theta,g}$ represents the contribution from the continuous spectrum.
+$[-\pi,\pi]_{per}$ denotes a periodic interval.
+The discrete spectrum is particularly important for insights into long-term behavior, such as periodicity and stability, while the continuous spectrum characterizes complex behaviors such as mixing, chaos, and transport phenomena [1, 47, 7].
+Our analysis emphasizes finding these spectral components including both discrete and continuous part from KMD.
+Extended Dynamic Mode Decomposition (EDMD) [44] is a widely used data-driven method for approximating the Koopman operator and its spectral components in KMD.
+It selects a set of observables $\{\psi_{i}\}_{i=1}^{N_{K}}$ to form a subspace $V_{N_{K}}=\text{span}\{\psi_{i}\}_{i=1}^{N_{K}}$.
+The method constructs a finite-dimensional approximation of the Koopman operator by solving a least-square problem based on system snapshots and selected dictionary, allowing the computation of eigenvalues, eigenfunctions, and Koopman modes.
+While common dictionaries include polynomials, Fourier basis, and RBF functions, selecting an optimal dictionary remains system-dependent and challenging.
+Given data snapshots $\{(x_{i},y_{i})\}_{i=1}^{m}$ with $y_{i}=F(x_{i})$ and $\{x_{i}\}_{i=1}^{m}$ i.i.d.
+by $\mu$, two data matrices $\Psi_{X}$ and $\Psi_{Y}$ are formed by evaluating the dictionary on these data points:
+$\Psi_{X}\coloneqq\begin{bmatrix}\psi_{1}(x_{1})&\dots&\psi_{N_{K}}(x_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(x_{m})&\dots&\psi_{N_{K}}(x_{m})\end{bmatrix},\quad\Psi_{Y}\coloneqq\begin{bmatrix}\psi_{1}(y_{1})&\dots&\psi_{N_{K}}(y_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(y_{m})&\dots&\psi_{N_{K}}(y_{m})\end{bmatrix}.$
+EDMD computes the approximated Koopman matrix representation by $K=\Psi_{X}^{\dagger}\Psi_{Y}$, where $\Psi_{X}^{\dagger}$ is the pseudoinverse of $\Psi_{X}$.
+The eigenvalues of $K$ provide approximation of the Koopman operator’s discrete spectrum, and the Koopman eigenfunctions $\phi_{i}$ are approximated as $\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i}$, where $\mathbf{v}_{i}\in\mathbb{C}^{N_{K}}$ is the $i$-th eigenvector of $K$ and $\mathbf{\Psi}\coloneqq[\psi_{1},\dots,\psi_{N_{K}}]$ is a (row) vector of observables from the dictionary.
 3 Koopman Operator Learning
-While EDMD approximates the Koopman operator, it suffers from spectral pollution as increasing dictionary sizes introduce spurious eigenvalues, obscuring the system’s true dynamics.
-Residual Dynamic Mode Decomposition (ResDMD) [12] addresses this by filtering out spurious eigenvalues using spectral residuals but relies on precomputed eigenpairs, inheriting inaccuracies from methods like EDMD.
-To overcome these limitations, we propose Neural Network-ResDMD (NN-ResDMD), which directly approximates the Koopman operator and its spectrum by minimizing a ResDMD-specific loss function.
-Using a Feedforward Neural Network (FNN), NN-ResDMD also optimizes dictionary functions for the Koopman invariant subspace, eliminating the need for manual basis selection.
+EDMD directly approximates the Koopman operator and its spectral components from data but suffers from spectral pollution, where discretization of the infinite-dimensional operator introduces spurious eigenvalues that have no relation to the true spectrum.
+Residual Dynamic Mode Decomposition (ResDMD) [8] improves spectral accuracy by filtering out these polluted eigenvalues via spectral residual measurement.
+However, it only filters precomputed spectra and cannot fully discover the Koopman operator’s complete spectral information, a limitation known as the ‘spectral inclusion’ problem where parts of the true spectrum may be missed entirely.
+To address these issues, we propose Residual-based Koopman-learning Network (ResKoopNet), which minimizes spectral residual over eigenpairs while exploring the optimal dictionary space.
+This approach identifies a more precise and complete spectrum of the Koopman operator while ensuring all learned dictionary functions satisfy small spectral residual criteria.
+We note that while we employ a simple neural network as our optimizer in this work, the specific choice of optimization tool can be adapted based on the data structure and properties of the dynamical system being studied.
 3.1 ResDMD Review
-Now, suppose we have obtained an eigenpair $(\lambda,\phi)$ of $\mathcal{K}$ from EDMD or other methods [13, 14, 15, 16, 17, 18] where $\lambda\in\mathbb{C}$ and the eigenfunction $\phi$ is expanded in terms of dictionary functions, i.e., $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}v_{i}\in V_{N_{K}}$ for some $\mathbf{v}\in\mathbb{C}^{N_{K}}$, where $v_{i}$ represents weights of the span.
-Without loss of generality, we consider $\phi$ has been normalized, i.e., $\|\phi\|_{2}=1$.
-The accuracy of this eigenpair approximation in the ResDMD framework can be measured by computing its squared relative residual using the dictionary in the following way:
+Now, suppose we have an estimated eigenvalue-eigenfunction pair $(\lambda,\phi)$ of $\mathcal{K}$ where $\lambda\in\mathbb{C}$ and $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}v_{i}\in V_{N_{K}}$ is expressed in the dictionary functions.
+Assuming $\phi$ is normalized, i.e., $\|\phi\|_{\mu}=1$, the spectral residual that measures the accuracy of this eigenpair is defined as:
 […]
-$=\sum_{i,j=1}^{N_{K}}\bar{v}_{i}\left[\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\lambda\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\bar{\lambda}\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}+|\lambda|^{2}\langle\psi_{i},\psi_{j}\rangle_{\mu}\right]v_{j},$ (3.1)
-where $\bar{v}_{i},\bar{\lambda}$ denote the complex conjugate of $v_{i},\lambda$.
-This squared relative residual in (A.1) is the theoretical value that measures the distance between $\phi$ and the eigenspace associated with $\lambda$, especially under the assumption that $\lambda$ is in the discrete spectrum of $\mathcal{K}$.
-To approximate this residual in practice, we apply the Galerkin approximation [19], which states that as the number of data points $m$ increases, the following limits hold:
+where $\bar{v}_{i},\bar{\lambda}$ denote the complex conjugates of $v_{i},\lambda$.
+This spectral residual in Eq.
+(3.1) quantifies how far the estimated eigenpair deviates from the true spectrum.
+In practice, we can approximate this residual using data samples through Galerkin methods [5, 8]; more specifically, the following convergence results hold as the number of data points $m\to\infty$:
 $\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{X}^{*}\Psi_{X}\right]_{ij}$ $=\langle\psi_{i},\psi_{j}\rangle_{\mu},$ (3.2)
 $\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{X}^{*}\Psi_{Y}\right]_{ij}$ $=\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu},$
-$\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{Y}^{*}\Psi_{Y}\right]_{ij}$ $=\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}=\langle\psi_{i},\mathcal{K}^{*}\mathcal{K}\psi_{j}\rangle_{\mu},$
-where * denotes complex conjugate. Using this approximation, the squared relative residual from (A.1) is approximated as follows (see A.1 for more details):
-$\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}\left[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}-\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}\right]\mathbf{v}.$ (3.3)
-where (3.3), denoted as $\widehat{\textit{res}}(\lambda,\phi)^{2}$, represents the approximation of the theoretical value in (A.1).
-By definition in (A.1), the residual quantifies the deviation from the spectral property, measuring how far the estimated eigenpair is from the true spectrum.
-In practice, (3.3) is calculated for all precomputed eigenpairs, retaining those with residuals below a threshold.
-However, while residuals help filter and select valid eigenpairs, they do not improve the accuracy of eigenpair estimation.
-3.2 Neural Network-ResDMD
-General framework In this section, we present the Neural Network-ResDMD (NN-ResDMD) framework, designed to compute the eigenpairs of the Koopman operator directly using ResDMD-based spectral residuals, as illustrated in Figure 1.
-The method first determines the optimal dictionary functions by minimizing the total residual $J\coloneqq\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}$, over all computed eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{N_{K}}$.
-The spectral residual directly impacts the finite-dimensional projection of the Koopman operator and our method minimizes this residual to ensure the learned basis functions adequately capture the Koopman dynamics.
-This approach allows the construction of the Koopman operator matrix $\tilde{K}$ without relying on external methods or post-processing.
-Equation (3.5) enables NN-ResDMD to compute eigenpairs directly, improving accuracy compared to ResDMD, which relies on filtering precomputed results from other methods.
-In this framework, neural networks parameterize the dictionary functions $\mathbf{\Psi}(x;\theta)$, where $\theta$ represents the network parameters.
-By minimizing the spectral residual $J$, this approach directly optimizes the dictionary functions towards better approximation of Koopman spectral components, which ensures the learned operator captures the underlying spectral properties of the dynamical system.
-This is fundamentally different from traditional methods like EDMD, which focus on minimizing prediction errors in the observable space without explicitly considering spectral accuracy.
-The neural network architecture serves as a flexible function approximator, that allows the framework to adaptively learn the optimal dictionary that minimizes spectral residuals, thereby producing more accurate and reliable Koopman spectral decompositions.
-This spectral-oriented optimization improves the accuracy of eigenvalues approximations and enhances the quality of the computed eigenfunctions, which leads to better characterization of the system’s dynamic behavior.
-Figure 1: (Left) The classical ResDMD and (Right) the Neural Networks based ResDMD methods
-From Residual to NN This section explains how neural networks are integrated into the ResDMD framework.
-In ResDMD, the squared relative residual approximation (3.3) measures how well a computed eigenpair fits the dataset.
-If the Koopman matrix $K$ is well-approximated by the projected Koopman operator $\mathcal{K}_{N_{K}}$, the total residual $J$ should approach zero as more data is provided.
-Thus, $J$ can be used as a loss function, and the optimal Koopman matrix $\tilde{K}$ is obtained by minimizing:
-$J=\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}.$ (3.4)
-which is equivalent to minimization the following (See A.2 for more details):
-$J=\frac{1}{m}\|(\Psi_{Y}-\Psi_{X}K)V\|_{F}^{2}$ (3.5)
-where $V$ is a matrix in which each column is an eigenvector $\mathbf{v}_{i}$ of Koopman matrix $K$.
-Thus, with a fixed dictionary function $\mathbf{\Psi}$, the explicit form for the optimal Koopman matrix $\tilde{K}$ can be directly computed as
-$\tilde{K}=G^{\dagger}A$ (3.6)
-where $G=\frac{1}{m}\Psi_{X}^{*}\Psi_{X},A=\frac{1}{m}\Psi_{X}^{*}\Psi_{Y}$.
-Remark.
-Typically a regularization term is needed to enhance stability. Here we add a small perturbation, i.e., $\tilde{K}=(G+\sigma I)^{-1}A$ for some small number $\sigma>0$.
-As shown in (3.6), NN-ResDMD provides an explicit expression for $\tilde{K}$ given the optimal dictionary function $\mathbf{\Psi}$, allowing for the direct computation of Koopman eigenpairs.
-The optimization problem in Equation 3.5 is to minimize the error along the eigen-basis, in contrast to the optimization problem $\|\Psi_{Y}-\Psi_{X}K\|_{F}^{2}$ for EDMD, thereby yielding different optimal $\mathbf{\Psi}$ compared to EDMD.
-Therefore, although the $K$ update procedure appear identical to the EDMD approach, they originate from different theoretical foundations and serve different optimization purposes.
-Additionally, it automatically optimizes basis functions using neural networks, removing the need for manual selection.
-Since NN-ResDMD is based on the ResDMD framework, it also retains the theoretical convergence guarantees that EDMD lacks: EDMD has convergence results under strong assumptions, such as requiring the Koopman operator to be bounded (Assumption 2 in [20]), ResDMD requires only that the operator is closed and densely defined.
-In NN-ResDMD, neural networks parameterize the dictionary functions $\mathbf{\Psi}(x;\theta)$ to minimize the total residual $J(\theta)$, as defined in (3.4).
-The feedforward neural network generates the dictionary functions based on data snapshots, and the total residual is given by:
-$J(\theta)=\frac{1}{m}\|(\Psi_{Y}(\theta)-\Psi_{X}(\theta)K(\theta))V(\theta)\|_{F}^{2}$ (3.7)
-where $K(\theta)$ and $V(\theta)$ depend on $\theta$. The Koopman matrix $\tilde{K}(\theta)$ is computed as:
-$\tilde{K}(\theta)=G(\theta)^{\dagger}A(\theta)$ (3.8)
-with $G(\theta)=\frac{1}{m}\Psi_{X}(\theta)^{*}\Psi_{X}(\theta)$ and $A(\theta)=\frac{1}{m}\Psi_{X}(\theta)^{*}\Psi_{Y}(\theta)$.
-The algorithm alternates between updating $K(\theta)$ via least squares and optimizing $\theta$ using gradient descent until $J(\theta)$ converges, yielding the approximated Koopman spectrum and optimized dictionary functions.
-While it is possible to optimize both $K(\theta)$ and $J(\theta)$ simultaneously, as done in [21] and [22], our separate procedure ensures computational efficiency and numerical stability compared to the coupled optimization case.
-Computing Algorithm In our neural networks implementation, we include some non-trainable basis outputs to enhance the dictionary functions.
-Specifically, we add a vector of ones and the coordinates of the state space as non-trainable basis in the output layer, which help avoid trivial solutions, i.e., $J=0$ for some initial $\theta$.
-For the network architecture, we build a three-layer Feedforward Network where each hidden layer size can be specified during training.
-We use the hyperbolic tangent (tanh) function as the activation function for the hidden layers.
-In terms of optimization, we employ the Adam optimizer for updating the network parameters.
-Adam is particularly well-suited for this task due to its ability to adapt the learning rate for each parameter, which can lead to faster convergence in the alternating optimization process between the network parameters and the Koopman matrix.
+$\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{Y}^{*}\Psi_{Y}\right]_{ij}$ $=\langle\psi_{i},\mathcal{K}^{*}\mathcal{K}\psi_{j}\rangle_{\mu},$
+where * denotes conjugate transpose of matrix or adjoint operator. Using these relationships, the spectral residual in Eq. (3.1) can be approximated as (see A.2 for more details):
+$\begin{split}\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}\\ -\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}]\mathbf{v}.\end{split}$ (3.3)
+Notice that Eq.
+(3.3) is calculated only for precomputed eigenpairs.
+However, the spectral residual serving as a threshold only filters polluted eigenpairs: it does not explore additional accurate eigenpairs or enhance the approximation.
+For an extreme case, the spectrum of the approximated Koopman operator could be trivial [8, Example 3.1].
+This is precisely the spectral inclusion problem we mentioned earlier, where the computed spectrum is constrained to a subset of the true spectrum, but could potentially omit critical spectral components.
+Addressing this fundamental limitation is a core objective of our work.
+Continuous Spectra and Pseudospectrum The concept of pseudospectrum is particularly useful when studying the continuous spectrum of the Koopman operator, as it captures regions where the resolvent norm (i.e., $\|(\mathcal{K}-\lambda I)^{-1}\|$) is large, even in the absence of discrete eigenvalues.
+In order to compute the pseudospectrum, a set of grid points as candidate spectrum values $\{z_{1},\ldots z_{n_{z}}\}$ is scanned in the complex plane using the spectral residual; specifically, for each grid point $z_{j}\in\mathbb{C}$, we compute the following SVD problem
+$\tau_{j}=\min_{\mathbf{v}_{i}\in\mathbb{C}^{N_{k}}}\widehat{\textit{res}}(z_{j},\mathbf{\Psi}\mathbf{v}_{i}),$ (3.4)
+which uses the dictionary $\mathbf{\Psi}$ computed from our proposed ResKoopNet method.
+The approximated $\epsilon$-pseudospectrum containing the continuous spectra is then given by $\{z_{j}:\tau_{j}<\varepsilon\}$.
+More details on proof of convergence can be found in Colbrook and Townsend [8, Appendix B].
+3.2 Residual-based Koopman-learning Network (ResKoopNet)
+General framework In this section, we present the Residual-based Koopman-learning Network (ResKoopNet), which minimizes spectral residual to address the spectral inclusion problem and yield a more accurate and complete Koopman spectrum.
+The overall pipeline is illustrated in Figure 1.
+ResKoopNet optimizes the dictionary functions by minimizing the total spectral residual,
+$J\coloneqq\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2},$
+over all computed eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{N_{K}}$.
+This loss directly influences the quality of the finite-dimensional Koopman approximation and ensures the learned dictionary captures essential spectral features without relying on precomputed spectra or post-processing.
+In this work, we parameterize the dictionary functions $\mathbf{\Psi}(x;\theta)$ using a simple feedforward neural network, as shown on the right side of Figure 1, where $\theta$ denotes the network parameters.
+Unlike EDMD, which minimizes prediction error, our method explicitly targets spectral accuracy.
+Remark 3.1.
+Notice that the neural network architecture can be replaced by more advanced models or other optimizers depending on the system and data structure.
+Figure 1: The optimizer on the right looks for the optimal dictionary by minimizing Eq. (3.3).
+In ResDMD’s framework, the spectral residual in Eq.
+(3.3) measures how well a computed eigenpair from the dataset, in other words, ideally we want the total spectral residual $J=\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}$ approaches to zero, which implies that we can treat $J$ as a loss function.
+Moreover, minimizing $J$ is also equivalent to the following minimization problem:
+$\min_{\tilde{K}}J=\min_{\tilde{K}}\frac{1}{m}\|(\Psi_{Y}-\Psi_{X}\tilde{K})V\|_{F}^{2},$ (3.5)
+where each column of the matrix $V$ is an (right) eigenvector $\mathbf{v}_{i}$ of the matrix $\tilde{K}$.
+Thus, with dictionary $\mathbf{\Psi}$ fixed, we can obtain a closed form for the optimal Koopman matrix $\tilde{K}$ as following
+$\tilde{K}=G^{\dagger}A,$ (3.6)
+where $G=\frac{1}{m}\Psi_{X}^{*}\Psi_{X},A=\frac{1}{m}\Psi_{X}^{*}\Psi_{Y}$ (See A.3 for more details on deriving the Eq. (3.5) and Eq. (3.6)).
+Although Eq.
+(3.6) resembles the EDMD formulation for computing the Koopman matrix representation $\tilde{K}$, ResKoopNet and ResDMD fundamentally differ from EDMD in their theoretical foundation.
+While they share a similar matrix expression, this expression is derived from minimizing the spectral residual in Eq.
+(3.3) rather than the prediction error that EDMD minimizes, as shown in Appendix A.3.
+Specifically, our approach explicitly incorporates $\mathcal{K}^{*}\mathcal{K}$ as shown in Eq.
+(3.2), focusing on spectral accuracy rather than observable prediction.
+This distinction leads to significantly different performance in capturing the spectral properties of the Koopman operator.
+Remark 3.2.
+To ensure the numerical stability when computing $G^{\dagger}$, we add a small perturbation, i.e., $\tilde{K}=(G+\sigma I)^{-1}A$ for some small number $\sigma>0$.
+As shown in Eq.
+(3.6), ResKoopNet provides an explicit expression for the optimal Koopman matrix $\tilde{K}$ given a fixed dictionary function $\mathbf{\Psi}$.
+Although the formulation is similar to EDMD, the key distinction between ResKoopNet and EDMD lies in their loss functions; more specifically, while EDMD minimizes the overall prediction error $\|\Psi_{Y}-\Psi_{X}\tilde{K}\|_{F}^{2}$, ResKoopNet minimizes the residual projected onto the estimated eigenspace, i.e., $\|(\Psi_{Y}-\Psi_{X}\tilde{K})V\|_{F}^{2}$.
+This eigenspace projection fundamentally changes the optimization landscape by weighting errors along spectral directions, ensuring that the learned Koopman matrix better captures the underlying eigenstructure rather than merely minimizing average prediction error.
+Therefore, they arise from distinct theoretical foundation.
+From Residual to Machine Learning This section explains how a simple neural network is integrated into ResKoopNet.
+In the optimizer as shown in Figure 1, the neural network parameterizes the dictionary $\mathbf{\Psi}(x;\theta)$ and minimize the parametrized total spectral residual $J(\theta)$ given in the following:
+$J(\theta)=\frac{1}{m}\|(\Psi_{Y}(\theta)-\Psi_{X}(\theta)K(\theta))V(\theta)\|_{F}^{2},$ (3.7)
+where $K(\theta)$ and $V(\theta)$ depend on the parameters $\theta$. The optimal Koopman matrix $\tilde{K}(\theta)$ is computed following Eq. (3.6):
+$\tilde{K}(\theta)=G(\theta)^{\dagger}A(\theta).$ (3.8)
+The algorithm alternates between updating $K(\theta)$ via Eq.
+(3.8) and optimizing $J(\theta)$ w.r.t.
+$\theta$ by stochastic gradient descent via Eq.
+(3.7).
+While it is possible to optimize both $K(\theta)$ and $J(\theta)$ simultaneously, as done in Takeishi et al.
+[35] and Otto and Rowley [30], our separate procedure ensures computational efficiency and numerical stability [23].
+Computing Algorithm In our neural network implementation, we include some non-trainable functions to enhance the dictionary computation.
+Specifically, we add a vector of constant one and the coordinates of the state space as non-trainable basis in the output layer, which helps avoid trivial solutions, i.e., $J=0$ for some initial $\theta$.
+For the network architecture, we build a simple feedforward neural network with three hidden layers, where each hidden layer size can be specified during training.
+We use the hyperbolic tangent (tanh) function as the activation function for the hidden layers and employ the Adam optimizer for updating the network parameters $\theta$.
+Adam is particularly well-suited for this task due to its ability to adapt the learning rate for each parameter, which can lead to faster convergence in the alternating optimization process between the network parameters $\theta$ and the Koopman matrix $K(\theta)$ formulation.
 The computing steps are illustrated in the following Algorithm 1.
 Input: Dataset $X,Y$, number of observables $N_{K}$, learning step $\delta$, regularization parameter $\sigma$, loss function threshold $\epsilon>0$, grid points $\{z_{1},\ldots z_{n_{z}}\}$.
-1 Initialize $\theta$, thus initializing $\mathbf{\Psi}(\theta)$ ;
+1 Initialize Initialize $\theta$, thus initializing $\mathbf{\Psi}(\theta)$ ;
 2 Compute $\tilde{K}(\theta)$ and its eigenvector matrix $V(\theta)$ ;
@@
 7
-8
-Output: $\tilde{K}(\theta)$, eigenpairs $\{(\lambda_{i},\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i})\}_{i=1}^{N_{K}}$ and pseudospectrum $\{z_{j}:\tau_{j}<\varepsilon\}$.
-9
+Output: $\tilde{K}(\theta)$, eigenpairs $\{(\lambda_{i},\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i})\}_{i=1}^{N_{K}}$ and pseudospectrum $\{z_{j}:\tau_{j}<\varepsilon\}$ (See Eq. (3.4)).
 Algorithm 1 NN-ResDMD
-While the practical advantages of NN-ResDMD are demonstrated through experiments, it is important to note its computational demands.
-The algorithm’s computational complexity stems primarily from its iterative optimization process.
-Each iteration involves a gradient descent update with complexity scaling linearly with both system dimensionality and neural network parameters.
-Although individual gradient steps are computationally lightweight for standard network architectures, the algorithm’s efficiency issue lies in its repeated least-squares optimizations.
-Compared to standard single least-squares computation as in most numerical algorithms, NN-ResDMD requires multiple iterations to achieve convergence, with stochastic gradient descent methods showing a theoretical $O(1/n)$ convergence rate.
-However, the method’s nonlinear optimization nature also presents challenges for establishing concrete convergence bounds and error estimates.
-If the continuous spectrum of the Koopman operator is of interest, following the ResDMD paper’s idea, we can scan candidate spectrum values within a grid in the complex plane using the residuals.
-Specifically, we compute $\tau_{j}=\min_{\mathbf{v}_{i}\in\mathbb{C}^{N_{k}}}\widehat{\textit{res}}(z_{j},\mathbf{\Psi}(\theta)\mathbf{v}_{i})$, where $\tau_{j}$ is the minimum residual for a grid point $z_{j}\in\mathbb{C}$.
-The approximated whole spectrum containing the continuous spectrum is then given by $\{z_{j}:\tau_{j}<\varepsilon\}$.
-More details can be found in [12].
-While the practical advantages of NN-ResDMD are demonstrated through experiments, it’s also worth noting that the method has theoretical underpinnings [23, 24] that support its convergence properties.
-A brief discussion on the convergence aspects of NN-ResDMD, leveraging existing results from approximation theory in Barron spaces, is provided in Appendix A.3.
-This discussion offers insights into how the neural network component of NN-ResDMD contributes to its effectiveness in approximating complex dynamical systems.
+ResKoopNet’s practical advantages come with computational demands due to its iterative optimization.
+Each gradient update scales linearly with system dimensionality and network parameters.
+While individual steps are lightweight, repeated least-squares optimizations make convergence slower than standard numerical methods, with stochastic gradient descent achieving an $O(1/n)$ rate.
+We also give a brief discussion on the convergence analysis of ResKoopNet in Appendix A.4, which leverages existing results from approximation theory in the Barron space [14, 42].
 4 Application in physical and biological systems
-In this chapter, we present three examples demonstrating the effectiveness of NN-ResDMD in estimating the key quantities of Koopman Mode Decomposition (KMD): spectrum, eigenfunctions, and Koopman modes.
-In the first low-dimensional example of a classical pendulum system, our method requires significantly fewer dictionary observables than [12, Section 4.3.1, Section 6.3] to compute the Koopman spectrum.
-The second high-dimensional example on turbulence highlights our method’s ability to detect acoustic vibrations and distinguish the pressure field through Koopman modes.
-The third example, a real-world high-dimensional neural system, compares NN-ResDMD with three popular methods—Hankel-DMD [25], EDMD with RBF basis, and kernelized-ResDMD [12]—and demonstrates its superiority in identifying and clustering latent dynamic structures.
-Together, these examples showcase NN-ResDMD’s performance across diverse systems and comprehensively evaluate its capabilities.
-Specifically, in all three experiments, we compare NN-ResDMD with Hankel-DMD, a theoretically grounded method that analyzes dynamical systems using time-delayed state measurements (see Appendix A.6.1 for a justification).
-Although its performance rivals NN-ResDMD in the simple pendulum system, it fails to capture key dynamics in higher-dimensional systems.
+In this section, we present three examples to demonstrate ResKoopNet’s effectiveness in estimating key components of Koopman Mode Decomposition: spectrum, eigenfunctions, and Koopman modes.
+In the pendulum example, we compare with EDMD, EDMD-DL, Hankel-DMD, and ResDMD, and show that ResKoopNet captures the full Koopman spectrum with significantly fewer dictionary observables than required by ResDMD, while other methods either produce spectral pollution or miss the continuous spectrum.
+Our turbulence analysis demonstrates ResKoopNet’s ability to detect not only acoustic vibration and turbulent fluctuation (as in ResDMD [8, Section 4.3.1, Section 6.3]) but also to recover the fundamental pressure field structure through the lowest-residual Koopman mode, which is a feature not captured by Hankel-DMD using the same dataset.
+In the neural dynamics example, we compare against three methods across five mice datasets, with ResKoopNet achieving superior clustering of neural states corresponding to different visual stimuli, as quantified by the Davies-Bouldin index.
+Note that we include Hankel-DMD [1] as a consistent benchmark across all experiments due to its theoretical guarantees and applicability to both low and high-dimensional systems without requiring predefined dictionaries.
+It uses time-delayed state measurements by Takens Embedding Theorem [36] instead of using a pre-defined dictionary or applying Galerkin approximation as in EDMD and ResDMD methods (see Appendix A.8.1 for more details.)
 4.1 Pendulum
-The pendulum system is a measure-preserving system due to its Hamiltonian nature, which theoretically implies its whole spectrum lies on the unit circle.
-For its dynamical behaviors, if the initial position of the pendulum is sufficiently far from the peak and the initial angular speed sufficiently small, the pendulum will oscillate; otherwise, the pendulum will pass the peak and rotate.
-In other words, this complex system exhibits two types of dynamical behaviors: rotation and oscillation.
-Here we simulate two cases with different numbers of initial points.
-We choose 90 and 240 initial points uniformly in the domain $[-\pi,\pi]_{per}\times[-15,15]$.
-Each point evolves 1000 steps with a step size of 0.5.
-Thus, the total data size in each set is approximately $9\times 10^{4}$ and $2.4\times 10^{5}$, respectively.
-As shown in Figure 4, only $N_{K}=300$ observables are needed to approximate the full spectrum, significantly fewer than the nearly 1000 required in [12, Section 4.3.1] for the same data size.
-Even with a much larger data size (Figure 4), the required observables remain small ($N_{K}=350$), demonstrating the robustness of efficient observables across data sizes.
-As shown in Figure 4, we compare our method to four approaches: EDMD, EDMD with Dictionary Learning (EDMD-DL), Hankel-DMD, and ResDMD, on the dataset with 90 initial points to compute the Koopman matrix and its corresponding spectral information.
-The first three methods (EDMD [11], EDMD-DL [18], and Hankel-DMD are limited to computing eigenvalues associated with the point spectrum.
-In these experiments, both EDMD and ResDMD use the hyperbolic cross approximation with Hermite functions up to order 15 and Fourier functions up to order 20.
-Hankel-DMD uses a time delay of 150.
-While Hankel-DMD yields accurate eigenvalues and shows points near the unit circle which matches the ground truth for this Hamiltonian system, it only captures the point spectrum(eigenvalues) and misses the entire spectrum(eigenvalues + continuous spectrum).
-It also still suffers from spectral pollution and requires careful tuning of the time delay parameter.
-With 300 basis functions, ResDMD is still unable to fully capture the entire spectrum, i.e., the unit circle, due to the insufficient number of basis functions.
-In the original ResDMD work, 964 basis functions using a hyperbolic cross approximation of order 100 were required to adequately cover the spectrum with a dataset of the same size [12, Section 4.3.1].
-Our method, in contrast, represents the results using shaded areas that show the pseudospectrum, which is a key feature that can capture the whole spectrum.
-While the shaded area may appear broad, this actually demonstrates our method’s ability to detect the complete spectrum.
-This radius of the shaded region accounts for computational uncertainties, as exact spectrum computation is computationally impossible.
-Theoretically, ResDMD guarantees that as this error tolerance(radius of the shaded region) approaches zero, the pseudospectrum converges to the true spectrum (the unit circle in this case) without spectral pollution.
-This comparison demonstrates that NN-ResDMD, even with only 300 basis functions, outperforms all four classical methods in terms of capturing the complete spectrum with greater accuracy and fewer basis functions.
-NN-ResDMD: $N_{K}=25$
-NN-ResDMD: $N_{K}=50$
-NN-ResDMD: $N_{K}=100$
-NN-ResDMD: $N_{K}=300$
-Figure 2: The four plots depict the spectrum of the Koopman operator, constructed using varying dictionary size $N_{K}$ of 25, 50, 100, and 300.
+The pendulum system is a measure-preserving Hamiltonian system, which theoretically constrains its spectrum to the unit circle. The system is governed by:
+$\ddot{\theta}=\sin(\theta),\quad(\theta,\dot{\theta})\in[-\pi,\pi]_{\text{per}}\times[-15,15],$
+where $\theta$ represents the angular displacement from equilibrium and $\dot{\theta}$ the angular velocity.
+The dynamics exhibit two distinct regimes: oscillatory motion for low-energy initial conditions, and rotational motion when the initial energy exceeds the separatrix threshold [24].
+For our analysis, we sample 90 and 240 initial conditions uniformly distributed in the domain, each evolving for 1000 time steps with $\Delta t=0.5$.
+This gives datasets of approximately $9\times 10^{4}$ and $2.4\times 10^{5}$ snapshots, respectively.
+We choose 3 hidden layers and use 300 neurons and 350 neuron in each hidden layer in both cases, respectively.
+[Uncaptioned image][Uncaptioned image]
+[Uncaptioned image][Uncaptioned image]
+Figure 2: The four plots depict the spectrum of the Koopman operator, constructed using varying dictionary sizes $N_{K}$ of 25, 50, 100, and 300.
 Each plot utilizes 90 initial points to illustrate the impact of increasing the dictionary size on approximating the spectrum of the Koopman operator.
-NN-ResDMD: $N_{K}=50$
-NN-ResDMD: $N_{K}=150$
-NN-ResDMD: $N_{K}=250$
-NN-ResDMD: $N_{K}=350$
-Figure 3: Same example as Figure 4 but with larger data size, using 240 initial points to show the effect of increasing dictionary size on approximating the Koopman operator spectrum.
+[Uncaptioned image][Uncaptioned image]
+[Uncaptioned image][Uncaptioned image]
+Figure 3: Same example as Figure 2 but with larger data size, using 240 initial points to show the effect of increasing dictionary size on the Koopman spectrum approximation.
 EDMD: $N_{K}=300$
@@
 ResDMD: $N_{K}=300$
-Figure 4: Comparison with classical methods.
-The four plots above represent the spectral information obtained from a $300\times 300$ Koopman matrix, calculated using four methods: EDMD, EDMD with Dictionary Learning (EDMD-DL), Hankel-DMD, and ResDMD.
-The illustrated eigenvalue spectra of the Koopman operator highlight the differences in results produced by these methods.
+Figure 4: Comparison with classical methods. Eigenvalue spectra computed from a $300\times 300$ Koopman matrix using EDMD, EDMD-DL, Hankel-DMD, and ResDMD.
+As shown in Figure 2, ResKoopNet requires only $N_{K}=300$ observables to approximate the full spectrum.
+While the original ResDMD work reported using 964 basis functions [8, Section 4.3.1], our additional tests indicate that approximately 460 basis functions can produce reasonable circular approximations with ResDMD (see Figure 7 in Appendix A.7.1).
+Nevertheless, ResKoopNet still achieves comparable accuracy with significantly fewer observables.
+This efficiency persists even with larger sizes of data points (Figure 3), where $N_{K}=350$ observables remain sufficient.
+Figure 4 compares ResKoopNet with four classical methods using the 90-initial-point dataset.
+Both EDMD and ResDMD employ Hermite functions (order 15) for $\theta$ and Fourier functions (order 20) for $\dot{\theta}$, while Hankel-DMD uses a time delay of 150.
+Though Hankel-DMD produces eigenvalues near the unit circle, it only captures the discrete point spectrum and cannot reconstruct the continuous spectral components.
+We include Hankel-DMD here for consistency in benchmarking across all experiments, despite its limitations in higher-dimensional systems examined later.
+The shaded region in the ResDMD plot represents the $\epsilon$-pseudospectrum, which approximates regions where the spectral residual is smaller than a threshold value; however, with 300 basis functions, ResDMD fails to fully recover the unit circle.
+Overall, this pendulum example demonstrates that ResKoopNet achieves accurate and complete spectral recovery using much fewer observables than other classical methods.
 4.2 Turbulence
-Recovering spatial patterns is a typical goal of DMD-based methods, especially in fluid dynamics, where Kernel ResDMD has been particularly successful in capturing such patterns and detecting acoustic vibrations in the turbulence system.
-However, Kernel ResDMD requires careful selection of kernel functions, while NN-ResDMD bypasses this by using neural networks to train observables and compute Koopman modes.
-We demonstrate this by applying NN-ResDMD to the turbulence system using the dataset from [12, Section 6.3].
-The ground truth in the first plot of Figure 5 represents a high-dimensional pressure field distribution (approximately 30,000 spatial dimensions) around an airfoil, with a clear distinction between the upper and lower surfaces.
-Technically, we apply truncated Singular Value Decomposition (SVD), select 300 observables, compute Koopman modes, and project them back into the original state space.
-In Figure 5, the first Koopman mode estimated by NN-ResDMD with the smallest residual value successfully highlights a clear global spatial separation that aligns with the pattern observed in the original pressure field.
-This advantage allows the first Koopman mode to directly distinguish spatial features present in the true pressure field, making it a powerful tool for interpreting complex fluid dynamics data.
-Subsequent Koopman modes also reveal strong acoustic waves that are critical in various aeronautical engineering fields.
-In contrast, Kernel ResDMD with a generic normalized Gaussian kernel function, as shown in the original work, is unable to produce a Koopman mode similar to the first Koopman mode from NN-ResDMD that clearly distinguishes the pressure field.
-For comparison, we also plot four Koopman modes computed by Hankel-DMD with a time delay of 5, corresponding to the four smallest residual values, which similarly do not reveal the pressure field patterns as in NN-ResDMD.
-These results are presented in Appendix Figure 7 and Appenxic A.6.2.
-Figure 5: The plots illustrate turbulence detection with Koopman modes computed by 300 observables.
-The first plot shows a 2D scatter plot of the pressure field, while the other plots display various Koopman modes, each labeled with corresponding residuals.
-The small residual values in the figures associated with the Koopman modes confirm the estimation accuracy.
-4.3 Identification of neural dynamics in mice visual cortex
-Since NN-ResDMD directly minimizes the residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics.
-To evaluate how effectively NN-ResDMD reveals latent temporal dynamics in real data, we apply it to a dataset of high-dimensional neural signals and demonstrate its advantages over a series of classical methods: the Hankel-DMD, EDMD (combined with RBF basis) and Kernel ResDMD.
+Recovering spatial patterns is a typical goal of DMD-based methods, especially in fluid dynamics [33, 28].
+Figure 5(a) shows the ground truth pressure distribution, where a clear asymmetry exists between the upper and lower airfoil surfaces.
+This fundamental spatial pattern is critical for analyzing aerodynamic performance and structural integrity.
+While the original Kernel-ResDMD method has successfully detected acoustic vibration and turbulent fluctuation shown in Colbrook and Townsend [8, Section 6.3], it fails to recover the global spatial structure of the pressure field.
+In contrast, ResKoopNet achieves a significant improvement by accurately reconstructing this dominant spatial pressure field structure itself from its leading Koopman mode as shown in Figure 5(b), which contains essential information about the flow characteristics around the airfoil.
+The turbulent flow dataset from Colbrook and Townsend [8, Section 6.3] models a two-dimensional airfoil system with Reynolds number $3.88\times 10^{5}$ and Mach number 0.07.
+The data captures a pressure field at 295,122 spatial points across 798 time steps, sampled every $2\times 10^{-5}$ seconds.
+Given the extreme dimensionality of this $X\in\mathbb{R}^{798\times 295,122}$ matrix, we apply truncated SVD on the data matrix to retain only the top 150 components and project the data via $XV=US\in\mathbb{R}^{798\times 150}$.
+After computing Koopman modes in this reduced state space using ResKoopNet, we map them back to the original dimensions through multiplication with $V$.
+To maintain consistency with the baseline, we use the same number of observables $N_{K}=250$ as in Colbrook and Townsend [8, Section 6.3], including 99 trainable neural network-based functions, one constant function, and 150 principal components obtained from the truncated SVD.
+Figure 5(b) shows the first Koopman mode computed by ResKoopNet, which corresponds to the mode whose eigenvalue has the largest absolute real part, and it also has the smallest spectral residual among all computed modes.
+This mode not only captures the most dominant dynamics in the system but also closely resembles the original pressure field.
+This correspondence demonstrates that ResKoopNet is able to recover the principal spatial feature of the turbulent flow, something that Kernel-ResDMD and Hankel-DMD could not achieve, which also justifies that the proposed method ResKoopNet can effectively address the spectral inclusion problem where traditional methods often fail to recover the complete spectrum of the Koopman operator.
+This mode likely corresponds to the largest singular value observed in the SVD, as shown in Figure 5(e) and (f).
+The dominant singular value highlights the strongest spatial mode in the data, which aligns with the physical air pressure field and its global structure.
+In addition to the first mode, Figure 5(c) and (d) show the third and seventh Koopman modes computed by ResKoopNet, associated with the eigenvalues having the third and seventh largest absolute values of their real parts.
+These two modes have the second and third smallest spectral residuals and correspond to acoustic vibration and turbulent fluctuation, as also observed in the Kernel-ResDMD study [8, Section 6.3].
+The ordering of the Koopman modes in Figure 5(b)–(d) is based on the ascending values of spectral residual, indicating how well each mode aligns with the spectrum of the underlying Koopman operator.
+For further comparison, we include in Appendix A.8.2 four Koopman modes computed using Hankel-DMD with a time delay of 5, selected similarly based on the smallest spectral residual.
+These modes, however, do not reveal the global pressure pattern observed in ResKoopNet’s first mode, which further validates the unique capability of ResKoopNet to recover large-scale coherent structures in turbulent systems.
+Refer to caption Refer to caption
+(a) (b)
+Refer to caption Refer to caption
+(c) (d)
+Refer to caption Refer to caption
+(e) (f)
+Figure 5: Turbulence detection using Koopman modes from 250 observables.
+(a) shows the original 2D pressure field.
+(b) shows Koopman mode 1 which has the smallest residual and closely matches the original pressure field.
+(c) and (d) show the acoustic vibration and turbulent fluctuation by the Koopman modes with the 2nd and 3rd smallest spectral residual.
+(e) and (f) show the plot of all singular values and first 150 singular values when applying truncated SVD method, respectively.
+4.3 Neural dynamics identification in mice visual cortex
+Figure 6: ResKoopNet outperforms Hankel-DMD in identifying latent dynamic structures in neural signals with a dictionary size of 50.
+(A) (Top) 50 Koopman eigenfunctions estimated by ResKoopNet across 6 states characterized by different video stimuli in an example mouse.
+Each trial contains 300 data points (10s at 50Hz).
+(Bottom) 50 Koopman eigenfunctions approximated by Hankel-DMD, each 50 points long, reflecting the dimension of the Hankel matrix.
+(B) 2D representation of Koopman eigenfunctions for all tested mice, computed by ResKoopNet and reduced via Multidimensional Scaling (MDS).
+Trials of the same state cluster well.
+(C) Same as (B) but computed with Hankel-DMD, showing no clear state separation.
+(D) 2D representation of Koopman eigenfunctions for the first mouse, computed by EDMD with an RBF basis.
+See Appendix Figure 14 for full results.
+(E) Same as (D) but computed with Kernel-ResDMD.
+See Appendix Figure 15 for full results.
+(F) Davies-Bouldin Indices (DBIs) evaluating clustering quality across four methods (ResKoopNet, Hankel-DMD, EDMD+RBF, and Kernel-ResDMD) for five mice.
+Lower DBI values for ResKoopNet indicate better clustering.
+Since ResKoopNet directly minimizes the residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics.
+To evaluate how effectively ResKoopNet reveals latent temporal dynamics in real data, we apply it to a dataset of high-dimensional neural signals and demonstrate its advantages over a series of classical methods: the Hankel-DMD, EDMD (combined with RBF basis) and Kernel-ResDMD.
 These methods are selected as representative approaches for handling high-dimensional data.
-The dataset is part of the open dataset on mice from the competition "Sensorium 2023" [26, 27].
+The dataset is part of the open dataset on mice from the competition “Sensorium 2023" [39, 40].
 In the experiments, mice viewed natural videos while their neural signals were recorded via calcium imaging in the primary visual cortex, reflecting the activity of thousands of neurons.
@@
 Each recording captures the activity of over 7,000 neurons, with each 10-second video sampled at 50 Hz, resulting in 300 data points per trial.
-We applied NN-ResDMD and three classical Koopman decomposition methods (Hankel-DMD, EDMD with RBF basis, and Kernel ResDMD) to these datasets, using different implementations and Koopman subspace dimensions.
-For NN-ResDMD, we trained dictionaries on all snapshots from each mouse to avoid overfitting, reduced the data to 300 dimensions via SVD, and selected 501 eigenfunctions.
+We applied ResKoopNet and three classical Koopman decomposition methods (Hankel-DMD, EDMD with RBF basis, and Kernel-ResDMD) to this dataset, using different implementations and Koopman subspace dimensions.
+For ResKoopNet, we trained dictionaries on all snapshots from each mouse to avoid overfitting.
+We reduced the data to 24 dimensions via truncated SVD, together with the constant and 25 trainable bases, resulting in 50 eigenfunctions.
+The neural network we use here consists of 3 hidden layers, each with 200 neurons.
+The hyperparameter scanning results are included in Appendix Figure 17, demonstrating that our parameter choice is within the robust range.
 The decomposed eigenfunctions are shown in Figure 6A(top), with markers indicating ground truth state separations.
@@
 In EDMD with RBF basis, we used the SVD-truncated 300 basis and 1000 RBF functions, resulting in 1301 eigenfunctions.
-For Kernel ResDMD, we used normalized Gaussians as kernel functions, setting the Koopman subspace dimension to 299 eigenfunctions based on [28].
+For Kernel-ResDMD, we used normalized Gaussians as kernel functions, setting the Koopman subspace dimension to 299 eigenfunctions based on Colbrook et al.
+[9].
 See Appendix A.9.4 for method details and Appendix A.10 for dictionary size justification.
-These eigenfunctions, shown in Figure 6A(bottom), Appendix Figure 9A, and Appendix Figure 10A, are compared to the ground truth trial identities.
+These eigenfunctions, shown in Figure 6A(bottom), Appendix Figure 14A, and Appendix Figure 15A, are compared to ground truth trial identities.
 The Koopman eigenfunctions represent dynamical features corresponding to the video stimuli.
-To evaluate their effectiveness, we assess how well eigenfunctions of the same stimuli cluster together, distinguishing them from other states.
+To evaluate their effectiveness, we assess how well eigenfunctions of the same stimuli cluster together, which distinguishes them from other states.
 If the eigenfunctions capture key dynamics related to the stimuli, those from trials with the same video should be separable from others.
 This turns the problem into a clustering task based on the separability of eigenfunctions across different stimuli.
-Note that averaged trial differences are even visibly clear for the NN-ResDMD case.
-Figure 6: NN-ResDMD outperforms Hankel-DMD in identifying latent dynamic structures in neural signals with a dictionary size of 501.
-(A) (Top) 500 Koopman eigenfunctions estimated by NN-ResDMD across 6 states characterized by different video stimuli in an example mouse.
-Each trial contains 300 data points (10s at 50Hz).
-(Bottom) 50 Koopman eigenfunctions approximated by Hankel-DMD, each 50 points long, reflecting the dimension of the Hankel matrix.
-(B) 2D representation of Koopman eigenfunctions for all tested mice, computed by NN-ResDMD and reduced via Multidimensional Scaling (MDS).
-Trials of the same state cluster well.
-(C) Same as (B) but computed with Hankel-DMD, showing no clear state separation.
-(D) 2D representation of Koopman eigenfunctions for the first mouse, computed by EDMD with an RBF basis.
-See Appendix Figure 9 for full results.
-(E) Same as (D) but computed with Kernel ResDMD.
-See Appendix Figure 10 for full results.
-(F) Davies-Bouldin Indices (DBIs) evaluating clustering quality across four methods (NN-ResDMD, Hankel-DMD, EDMD+RBF, and Kernel ResDMD) for five mice.
-Lower DBI values for NN-ResDMD indicate better clustering.
+Note that averaged trial differences are even visibly clear for the ResKoopNet case.
+It is worth clarifying that clustering these processes is not biologically necessary, as simple clustering of mean firing rates is sufficient (see Appendix Figures 1112).
+However, this dataset provides an ideal benchmark with clear ground truth labels, enabling accurate comparison of Koopman eigenfunction estimation methods.
+Effective estimation of Koopman eigenfunctions can be valuable for more complex tasks, such as unsupervised latent state identification or decoding object dynamics from video stimuli, offering promising directions for future research.
 We use Multi-dimensional Scaling (MDS) to visualize how these eigenfunction-based features cluster according to ground truth states.
 MDS reduces data dimensionality based on similarities, making it ideal for visualizing clustering performance.
-While UMAP and t-SNE are alternative methods, we show MDS results in 2D space (Figure 6B-E), with similar results for UMAP and t-SNE in the supplementary materials (Appendix Figure 8, Appendix Figure 9C,D and Appendix Figure 10C,D).
-The 2D MDS visualization reveals clear separation of features for all 5 mice using NN-ResDMD (Figure 6B), whereas no other method shows clear clustering (Figure 6C-E, Appendix Figure 9B, Appendix Figure 10B).
+While Uniform Manifold Approximation and Projection (UMAP) and t-distributed Stochastic Neighbor Embedding (t-SNE) are alternative methods, we show MDS results in 2D space (Figure 6B-E), with similar results for UMAP and t-SNE in Appendix Figure 13, Appendix Figure 14C, D and Appendix Figure 15C, D.
+The 2D MDS visualization reveals a clear separation of features for all 5 mice using ResKoopNet (Figure 6B), whereas no other method shows clear clustering (Figure 6C-E, Appendix Figure 14B, Appendix Figure 15B).
 To quantify this clustering, we calculate the Davies-Bouldin index (DBI), a measure of clustering quality that assesses how compact and well-separated the clusters are.
 A lower DBI indicates more compact clusters that are farther apart from each other, which corresponds to better clustering.
-The DBI is significantly lower for NN-ResDMD (Figure 6F), suggesting that it captures the latent dynamic structure more effectively than all three other methods.
+The DBI is significantly lower for ResKoopNet (Figure 6F), suggesting that it captures the latent dynamic structure more effectively than all three other methods.
 Similar clustering patterns are confirmed with UMAP and t-SNE (Appendix Figure 16).
 5 Conclusion and future work
-Koopman spectral components (eigenpairs) are fundamental to understanding dynamical systems, as they reveal intrinsic patterns and structures underlying complex temporal behavior through a linear framework for analyzing nonlinear dynamics.
-In this paper, we introduced NN-ResDMD, a method for estimating eigenpairs by minimizing spectral residuals, eliminating ResDMD’s need to filter precomputed results.
-Despite higher computational costs, using neural networks to learn eigenpairs provides a significant advantage by capturing patterns automatically and reducing manual intervention in basis selection.
-This flexibility is particularly beneficial for high-dimensional systems where traditional methods often struggle.
-Our experiments demonstrate that NN-ResDMD outperforms classical methods—including EDMD, Hankel-DMD, ResDMD, and their variants—in uncovering critical spatiotemporal characteristics of nonlinear dynamics.
-Despite the advantages, NN-ResDMD has several limitations and we discuss the major ones here.
-First, the neural network structure incurs higher computational costs compared to classical approaches, making it unsuitable for real-time learning tasks (see a brief discussion in Appendix A.5).
-Second, the deterministic nature of the framework does not account for stochastic aspects of the system, such as those addressed by methods like VAMP [29], limiting its applicability to highly noisy data.
-Additionally, the performance of NN-ResDMD is sensitive to hyperparameter tuning, including network architecture, dictionary size, and training criteria, which can require significant effort to optimize.
-Koopman eigenpairs provide unique perspectives into the interpretation of nonlinear dynamical mechanisms, and feedforward neural networks (FNNs) represent an initial step in learning spectral properties directly from data.
-In recent years, various deep neural network structures have been employed to learn the Koopman representations with different optimization targets other than the spectral residuals (e.g.
-[30, 21, 29, 31, 22, 32, 33, 34](see Appendix section A.4 for a comparison with the VAMP framework).
-With our approach, we demonstrate that even basic architectures can achieve significant improvements in Koopman operator estimation by using the spectral residual loss.
-Therefore, future work could focus on refining neural network architectures to enhance the accuracy and efficiency of Koopman eigenpair estimation.
-One promising direction is the incorporation of Physics-Informed Neural Networks (PINNs) and Physics-Informed Neural Operators (PINOs), which integrate physical laws directly into the learning process.
-This integration will ensure that the resulting Koopman eigenfunctions align with known physical constraint, avoid overfitting and faciliatates generalization.
-Indeed, the integration of PINNs and PINOs with the Koopman framework has the potential to serve as a powerful bridge between data-driven and model-driven approaches, offering enhanced insights into complex systems and enabling more robust temporal evolution predictions.
+In this paper, we introduced ResKoopNet, a novel approach that minimizes spectral residuals to accurately approximate Koopman operators for complex dynamical systems.
+This method successfully resolves the spectral inclusion problem while capturing both discrete and continuous spectra with high fidelity.
+Our experiments across physical and biological systems demonstrate that ResKoopNet significantly outperforms existing methods in identifying coherent structures and uncovering latent dynamics, particularly in high-dimensional applications.
+Despite these strengths, ResKoopNet has several limitations.
+First, the use of neural networks increases computational cost compared to classical algorithms (see Appendix A.6).
+Second, unlike stochastic approaches such as VAMP [25] and SDMD [46], ResKoopNet currently does not account for stochasticity (see Appendix A.5 for comparison).
+In addition, its performance can vary depending on the neural architecture and training configuration.
+Many physical laws, such as conservation principles or symmetry constraints, are inherently reflected in the spectral properties of dynamical systems.
+Incorporating such physical structure into the neural network architecture may therefore enhance the learning of Koopman spectral information.
+Future work could explore physics-informed designs that embed prior knowledge directly into the learning process, allowing the network to better align with the system’s underlying dynamics.
+This integration may improve both the interpretability and accuracy of the extracted Koopman representations.
 Acknowledgments
-Authors are grateful to Matthew Colbrook for valuable guidance on ResDMD technical part and providing fluid data, to Bin Han for discussion on computational issues, to Michel Beserve and Rory Bufacchi for proofreading and to Turishcheva Polina for providing permission of using the Sensorium 2023 dataset.
+Authors are grateful to Matthew Colbrook for valuable guidance on ResDMD technical part and providing fluid data, to Bin Han for discussion on computational issues, to Rouslan Krechetnikov for discussion on turbulence analysis, to Michel Beserve and Rory Bufacchi for proofreading and to Turishcheva Polina for providing permission of using the Sensorium 2023 dataset.
 Appendix A Appendix
-A.1 Calculation steps for 3.3
-Here we are going to show how squared relative residual implies (A.1) and then implies (3.3).
-Consider $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}\mathbf{v}_{i}$ with $\|\phi\|_{2}=1$, then
+A.1 Source Code
+For reproducibility, the source code will be available at the following link.
+A.2 Calculation steps for Eq. (3.3)
+Here we are going to show how spectral residual implies (3.3).
+Consider $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}\mathbf{v}_{i}$ with $\|\phi\|_{\mu}=1$, then
 […]
@@
 $=\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}\mathbf{v}_{j}-\bar{\lambda}\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}\mathbf{v}_{j}-\lambda\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}\mathbf{v}_{j}+|\lambda|^{2}\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\langle\psi_{i},\psi_{j}\rangle_{\mu}\mathbf{v}_{j}$
-[…]
+$=\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\left[\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\bar{\lambda}\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\lambda\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}+|\lambda|^{2}\langle\psi_{i},\psi_{j}\rangle_{\mu}\right]\mathbf{v}_{j}$
 $\approx\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\left[\frac{1}{m}[\Psi_{Y}^{*}\Psi_{Y}]_{ij}-\bar{\lambda}\frac{1}{m}[\Psi_{X}^{*}\Psi_{Y}]_{ij}-\lambda\frac{1}{m}[\Psi_{Y}^{*}\Psi_{X}]_{ij}+|\lambda|^{2}\frac{1}{m}[\Psi_{X}^{*}\Psi_{X}]_{ij}\right]\mathbf{v}_{j}$
 […]
-Remark.
-[…]
-A.2 Details for deriving (3.5)
+A.3 Details for deriving (3.5)
 $J$ $=\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}$
@@
 $=\sum_{i=1}^{N_{K}}\frac{1}{m}\langle\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i},\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i}\rangle$
-$=\sum_{i=1}^{N_{K}}\frac{1}{m}\|\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i}\|_{2}^{2}$
+$=\sum_{i=1}^{N_{K}}\frac{1}{m}\|\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i}\|_{\mu}^{2}$
 $=\frac{1}{m}\|(\Psi_{Y}-\Psi_{X}K)V\|_{F}^{2}.$
@@
 $0=\frac{dJ}{dK}$ $=\frac{d\operatorname{tr}(J)}{dK}\quad(\text{since $J$ is a scalar})$
-$=\frac{d}{dK}\operatorname{tr}\bigg{(}\frac{1}{m}\sum_{i=1}^{N_{K}}\mathbf{v}_{i}^{*}\bigg{[}\Psi_{Y}^{*}\Psi_{Y}-(\Psi_{X}^{*}\Psi_{Y})^{*}K$
-$\qquad-K^{*}(\Psi_{X}^{*}\Psi_{Y})+K^{*}(\Psi_{X}^{*}\Psi_{X})K\bigg{]}\mathbf{v}_{i}\bigg{)}$
+$=\frac{d}{dK}\operatorname{tr}\bigg{(}\frac{1}{m}\sum_{i=1}^{N_{K}}\mathbf{v}_{i}^{*}\bigg{[}\Psi_{Y}^{*}\Psi_{Y}-(\Psi_{X}^{*}\Psi_{Y})^{*}K-K^{*}(\Psi_{X}^{*}\Psi_{Y})+K^{*}(\Psi_{X}^{*}\Psi_{X})K\bigg{]}\mathbf{v}_{i}\bigg{)}$
 $=\sum_{i=1}^{N_{K}}\frac{d}{dK}\operatorname{tr}\bigg{(}\mathbf{v}_{i}^{*}\bigg{[}L-A^{*}K-K^{*}A+K^{*}GK\bigg{]}\mathbf{v}_{i}\bigg{)}$
@@
 $-2A+2GK=0\Rightarrow K=G^{\dagger}A.$
-Remark.
+Remark A.1.
 To solve $\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)$ , we simply rewrite it as
 $\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)=\tfrac{d}{dK}\operatorname{tr}\left((K\mathbf{v}_{i})^{*}G(K\mathbf{v}_{i})\right)$
-A.3 Discussion On Convergence
-To understand how neural networks enhance NN-ResDMD, it is important to introduce Barron space [35, 36, 23, 37].
-Barron space characterizes functions efficiently approximated by two-layer neural networks, which is central to NN-ResDMD.
-By leveraging networks that approximate functions within this space, NN-ResDMD can flexibly optimize the dictionary functions for Koopman operator approximation, making it highly effective for complex, high-dimensional systems.
+A.4 Discussion on convergence analysis
+To understand how neural networks enhance ResKoopNet and build up the theoretical framework of convergence, it is important to first introduce Barron space [31, 10, 14, 2].
+Barron space characterizes functions efficiently approximated by two-layer neural networks, which is central to ResKoopNet.
+By leveraging networks that approximate functions within this space, ResKoopNet can flexibly optimize the dictionary functions for Koopman operator approximation, making it highly effective for complex, high-dimensional systems.
 A function $f$ belongs to Barron space $\mathcal{B}$ if it can be represented as:
@@
 $\|f\|_{\mathcal{B}}=\inf_{\rho\in P_{f}}\left(\int_{\Omega}|a|\|w\|_{1}\rho(da,dw)\right),$
-where $P_{f}$ is the set of distributions for which $f$ can be represented. This framework provides a basis for analyzing approximation errors in neural networks.
+where $P_{f}$ is the set of distributions for which $f$ can be represented. This framework provides a foundation for analyzing approximation errors in neural networks.
 The following theorem [11] discusses the approximation capabilities of two-layer neural networks within this context, establishing a foundation for the subsequent analysis.
@@
 For any $f\in\mathcal{B}$ and $r\in\mathbb{N}$, there exists a two-layer neural network $f_{r}$ with $r$ neurons $\{(a_{i},\mathbf{w}_{i})\}$ such that
-$\|f-f_{r}\|_{L^{2}}\lesssim\frac{\|f\|_{\mathcal{B}}}{\sqrt{r}}.$
-This result implies that the approximation error decreases at a rate of $1/\sqrt{r}$ as the number of neurons $r$ increases, with the constant $\|f\|_{\mathcal{B}}$ reflecting the complexity of the function $f$ within the Barron space.
-Now, consider a Barron space $\mathcal{B}$ which is dense in $L^{2}(\Omega,\mu)$ and a projected Koopman operator $\mathcal{K}_{N_{K}}:\mathcal{B}_{N_{K}}\to L^{2}(\Omega,\mu)$ where $\mathcal{B}_{N_{K}}\subseteq\mathcal{B}$ is a $N_{K}$-dimensional subspace spanned by some dictionary $\mathbf{\Psi}=\{\psi_{i}\}_{i=1}^{N_{K}}$.
-According to Theorem A.1, we can have a well-trained dictionary that almost spans $\mathcal{B}_{N_{K}}$, i.e., given $\epsilon>0$, we can always obtain a dictionary $\mathbf{\Psi}_{r}=\{\psi_{r,i}\}_{i=1}^{N_{K}}$ such that $\sum_{i=1}^{N_{K}}\|\psi_{r,i}-\psi_{i}\|^{2}_{2}<\epsilon$.
-A.4 Highlights of NN-ResDMD compared with typical existing neural network-based Koopman framework
-Our NN-ResDMD method takes a fundamentally different approach from existing deep learning methods by building upon the residual-based framework of ResDMD rather than the different Koopman-approximating loss functions following the variational principles of VAMPnets [39, 40, 29] or the deep autoencoder structure in [41].
+$\|f-f_{r}\|_{\mu}\lesssim\frac{\|f\|_{\mathcal{B}}}{\sqrt{r}}.$
+This implies an approximation error decreasing at a rate of $O(1/\sqrt{r})$, where $r$ is the number of neurons.
+In ResKoopNet, the dictionary $\Psi(x;\theta)=\{\psi_{i}(x;\theta)\}_{i=1}^{N_{K}}$ is parameterized by a neural network with parameters $\theta$.
+Assuming the true dictionary functions $\psi_{i}\in\mathcal{B}$, Theorem A.2 ensures that $\Psi(x;\theta)$ can approximate the optimal dictionary spanning the Koopman invariant subspace $\mathcal{B}_{N_{K}}\subset\mathcal{F}$ with error $\epsilon>0$, provided $r$ is sufficiently large.
+Remark A.3.
+Notice that, the “two-layer neural network" in the Theorem A.2 statement refers to a hidder layer + an output layer, which is the most standard and general setting.
+Our implementation uses three hidden layers.
+This does not invalidate the result, as deeper networks can achieve at least the same approximation power [31].
+We want to show two convergence results here: (1) $\Psi(x;\theta)$ approaches the true invariant subspace of $\mathcal{K}$; (2) The eigenpairs $(\lambda_{i},\phi_{i})$ and pseudospectrum approximate $\mathcal{K}$’s true spectrum as $J(\theta)\to 0$.
+Assumption A.4.
+To formalize convergence, we make the following assumptions:
+  1. (a)
+    The optimal dictionary functions $\{\psi_{i}^{*}\}_{i=1}^{N_{K}}$ spanning $\mathcal{K}$’s invariant subspace lie in $\mathcal{B}$.
+  2. (b)
+    The loss $J(\theta)$ is Lipschitz continuous in $\theta$, and the neural network $\Psi(x;\theta)$ has bounded gradients, facilitating gradient-based optimization.
+Now, consider a Barron space $\mathcal{B}$ which is dense in $\mathcal{F}$.
+Given $N_{K}$ fixed, let $\mathcal{B}_{N_{K}}=\text{span}\{\psi_{i}^{*}\}_{i=1}^{N_{K}}$ be the true invariant subspace.
+By Theorem A.2, for each $\psi_{i}^{*}$, there exists a neural network approximation $\psi_{i}(x;\theta_{r})$ with $r$ neurons such that:
+$\|\psi_{i}^{*}-\psi_{i}(\cdot;\theta_{r})\|_{\mu}\leq\frac{\|\psi_{i}^{*}\|_{\mathcal{B}}}{\sqrt{r}}.$
+The total dictionary error is:
+$\|\Psi^{*}-\Psi(\cdot;\theta_{r})\|_{F}^{2}=\sum_{i=1}^{N_{K}}\|\psi_{i}^{*}-\psi_{i}(\cdot;\theta_{r})\|_{\mu}^{2}\leq\frac{1}{r}\sum_{i=1}^{N_{K}}\|\psi_{i}^{*}\|_{\mathcal{B}}^{2}=\frac{C_{N_{K}}}{r},$
+where $C_{N_{K}}=\sum_{i=1}^{N_{K}}\|\psi_{i}^{*}\|_{\mathcal{B}}^{2}$ is finite under Assumption A.4(a).
+As $r\to\infty$, $\Psi(x;\theta_{r})\to\Psi^{*}$ in the Frobenius norm, which ensures the dictionary approximated by neural network can represent $\mathcal{B}_{N_{K}}$.
+Algorithm 1 updates $\theta$ via stochastic gradient descent (SGD) with step size $\delta$ and computes $\tilde{K}(\theta_{n})$ iteratively until $J(\theta_{n})<\epsilon$ where $\theta_{n}$ represents $n$-th iteration of parameters $\theta$.
+For a Lipschitz continuous $J(\theta)$ with constant $L$ (by Assumption A.4(b)) and a strongly convex region near the optimum $\theta^{*}$ (assumed locally for simplicity), SGD converges at a rate of $O(1/n)$ in expectation [4]:
+$\mathbb{E}[J(\theta_{n})-J(\theta^{*})]\leq\frac{L}{2\eta n},$
+where $\eta$ is the strong convexity constant and $n$ is the iteration number.
+In practice, $J(\theta)$ is non-convex due to the neural network, and the alternating update with $\tilde{K}(\theta)$ stabilizes the process.
+As iteration step $n\to\infty$ and amount of data points $m\to\infty$, $J(\theta_{n})\to 0$, which implies $\widehat{\text{res}}(\lambda_{i},\phi_{i})\to 0$ for all $i$.
+With $J(\theta)\to 0$, we now assess the spectral error: if $\widehat{\text{res}}(\lambda_{i},\phi_{i})<\epsilon$, then $\|\mathcal{K}\phi_{i}-\lambda_{i}\phi_{i}\|_{\mu}/\|\phi_{i}\|_{\mu}<\sqrt{\epsilon}$, indicating $\lambda_{i}$ and $\phi_{i}$ are approximate eigenpairs of $\mathcal{K}$ converges to $\mathcal{K}$’s true spectrum as $N_{K}\to\infty$ and $\epsilon\to 0$, as pointed out in [8, Theorem B.1].
+Uniform convergence on compact subsets of $\mathbb{C}$ follows from the density of $\mathcal{B}_{N_{K}}$ in $\mathcal{F}$ and Dini’s theorem, as pointed out in Colbrook and Townsend [8, Lemma B.1].
+The non-convexity and system complexity may slow the spectral approximation in practice.
+High-order convergence (e.g., polynomial) could arise for smooth dynamics (See Colbrook and Townsend [8, Theorem 3.1] for more details), which is useful for further study.
+Remark A.5.
+The computational cost (Appendix A.6) scales with $r$, $t$, and $m$, which trades off with accuracy. Adaptive selection of $r$ and early stopping when $J(\theta)<\epsilon$ can optimize this balance.
+A.5 Highlights of ResKoopNet compared with typical existing neural network-based Koopman framework
+Our ResKoopNet method takes a fundamentally different approach from existing deep learning methods by building upon the residual-based framework of ResDMD rather than the different Koopman-approximating loss functions following the variational principles of VAMPnets [37, 45, 25] or the deep autoencoder structure in [24].
 By incorporating spectral residual measures into deep learning and introducing a structured representation that captures dependencies among eigenvalues, we achieve more compact and interpretable models for nonlinear systems with continuous spectra.
@@
 However, although they both depend on the covariance matrices (in our manuscript Equation 3.2), their methodologies differ significantly.
-Our residual-based method directly minimizes the spectral approximation error of the Koopman operator and accommodates both point and continuous spectra, while the VAMP score follows a variational framework, maximizing the sum of singular values to approximate the point spectrum, primarily for stochastic systems (though see an exception in TianWu+2021+635+659 [39]).
+Our residual-based method directly minimizes the spectral approximation error of the Koopman operator and accommodates both point and continuous spectra, while the VAMP score follows a variational framework, maximizing the sum of singular values to approximate the point spectrum, primarily for stochastic systems (though see an exception in Tian and Wu [37]).
 Moreover, while VAMP is specifically designed for Markov processes and requires the Koopman operator to be Hilbert-Schmidt, our approach focuses on deterministic systems and enables a more comprehensive spectral analysis that incorporates continuous spectra.
 This distinction in scope and methodology highlights how the two frameworks complement each other in addressing different aspects of spectral estimation.
-A.5 Discussion of computation costs
-Despite the various advantages of the NN-ResDMD framework, one significant limitation is its higher computational cost compared to the original ResDMD and other classical methods.
-Theoretical Perspective: The NN-ResDMD algorithm’s computational demands stem primarily from its iterative optimization process.
+A.6 Discussion of computation costs
+Despite the various advantages of the ResKoopNet framework, one significant limitation is its higher computational cost compared to the original ResDMD and other classical methods.
+Theoretical Perspective: The ResKoopNet algorithm’s computational demands stem primarily from its iterative optimization process.
 Each iteration involves a gradient descent update with complexity scaling linearly with both system dimensionality and neural network parameters.
@@
 Below we provide a detailed comparison between computation costs of several methods we used in this paper.
-Comparison over methods: The computational costs of EDMD, ResDMD, EDMD-DL, Hankel-DMD, and NN-ResDMD vary significantly based on their core computational steps and specific configurations.
+Comparison over methods: The computational costs of EDMD, ResDMD, EDMD-DL, Hankel-DMD, and ResKoopNet vary significantly based on their core computational steps and specific configurations.
 For a dataset with $m=10^{5}$ data points and $N_{K}=300$ dictionary functions (for EDMD-based methods), the theoretical complexity and runtime differ across methods.
@@
 This leads to runtimes ranging from minutes to hours, depending on the resolution of the pseudospectrum grid.
-EDMD-DL incorporates dictionary learning through stochastic gradient descent (SGD), where each iteration involves matrix construction ($O(N_{K}^{2}m)$), Koopman matrix computation ($O(N_{K}^{3})$), and neural network forward/backward propagation ($O(d|H|)$, with $d|H|$ representing the total network parameter size).
-With $k$ SGD iterations, the total complexity becomes $O(k(N_{K}^{2}m+N_{K}^{3}+d|H|))$, leading to runtimes also in the range of minutes to hours depending on $k$.
-NN-ResDMD, which builds on EDMD-DL, shares the same complexity, $O(k(N_{K}^{2}m+N_{K}^{3}+d|H|))$, but includes the explicit use of Koopman matrix eigenvectors and optional pseudospectrum computation, making its runtime slightly longer than EDMD-DL for high-resolution spectral analysis.
+EDMD-DL incorporates dictionary learning through stochastic gradient descent (SGD), where each iteration involves matrix construction ($O(N_{K}^{2}m)$), Koopman matrix computation ($O(N_{K}^{3})$), and neural network forward/backward propagation ($O(|\theta|)$, with $|\theta|$ representing the total network parameter size).
+With $k$ SGD iterations, the total complexity becomes $O(k(N_{K}^{2}m+N_{K}^{3}+|\theta|))$, leading to runtimes also in the range of minutes to hours depending on $k$.
+ResKoopNet, which builds on EDMD-DL, shares the same complexity, $O(k(N_{K}^{2}m+N_{K}^{3}+|\theta|))$, but includes the explicit use of Koopman matrix eigenvectors and optional pseudospectrum computation, making its runtime slightly longer than EDMD-DL for high-resolution spectral analysis.
 Hankel-DMD, using a time delay embedding dimension $T$, constructs a Hankel matrix ($O(Tm)$), performs singular value decomposition (SVD) ($O(T^{2}m)$), and computes the eigenvalues of a reduced $T\times T$ matrix ($O(T^{3})$).
 The total complexity is $O(Tm+T^{2}m+T^{3})$, and the runtime is heavily influenced by $T$, typically ranging from minutes to hours.
-While EDMD is computationally the most efficient, ResDMD and Hankel-DMD provide higher precision and robustness in spectral analysis at the expense of increased runtime, and EDMD-DL and NN-ResDMD offer flexibility and accuracy through dictionary learning, with additional SGD iterations and optional pseudospectrum computation contributing to their computational burden.
+While EDMD is computationally the most efficient, ResDMD and Hankel-DMD provide higher precision and robustness in spectral analysis at the expense of increased runtime, and EDMD-DL and ResKoopNet offer flexibility and accuracy through dictionary learning, with additional SGD iterations and optional pseudospectrum computation contributing to their computational burden.
 Empirical Perspective: In our experiments, without computing the pseudospectrum, the computational cost of ResDMD typically ranges from seconds to minutes.
-NN-ResDMD, on the other hand, can require tens of minutes to several hours, depending on factors such as data dimensionality, the number of snapshots, hidden layer configurations, dictionary sizes, and training convergence criteria.
-Trade-off Between Cost and Accuracy: While NN-ResDMD’s additional computational steps introduce higher costs, they enhance the accuracy and robustness of Koopman eigenpair estimation by allowing automatic dictionary learning and minimizing spurious spectral components.
-This trade-off makes NN-ResDMD particularly valuable in applications where precision is critical.
+ResKoopNet, on the other hand, can require tens of minutes to several hours, depending on factors such as data dimensionality, the number of snapshots, hidden layer configurations, dictionary sizes, and training convergence criteria.
+Trade-off Between Cost and Accuracy: While ResKoopNet’s additional computational steps introduce higher costs, they enhance the accuracy and robustness of Koopman eigenpair estimation by allowing automatic dictionary learning and minimizing spurious spectral components.
+This trade-off makes ResKoopNet particularly valuable in applications where precision is critical.
 However, its computational demands render it less suitable for real-time or online Koopman model learning tasks.
-A.6 Hankel-DMD
-A.6.1 Justification of Using Hankel-DMD as comparison in all experiments
+Detailed Analysis of Computational Bottlenecks: The complexity $O(k(N_{K}^{2}m+N_{K}^{3}+|\theta|))$ arises from ResKoopNet’s iterative optimization, alternating between updating the neural network parameters $\theta$ and computing the Koopman matrix $\hat{K}(\theta)$.
+The dominant term $O(kN_{K}^{2}m)$ reflects the cost of constructing data matrices $\Psi_{X}(\theta)$ and $\Psi_{Y}(\theta)$ at each iteration, followed by gradient computation.
+For high-dimensional systems (large $m$) or complex architectures (large $|\theta|$), the number of iterations $k$ required for convergence increases, particularly if the loss $J(\theta)=\frac{1}{m}\|(\Psi_{Y}(\theta)-\Psi_{X}(\theta)K(\theta))V(\theta)\|_{F}^{2}$ is non-convex or ill-conditioned.
+The term $O(kN_{K}^{3})$ stems from solving $\hat{K}(\theta)=(G(\theta)+\sigma I)^{-1}A(\theta)$, where matrix inversion scales cubically with $N_{K}$.
+The optional pseudospectrum computation, with complexity $O(n_{z}N_{K}^{3})$, becomes significant for systems with continuous spectra, as $n_{z}$ grid points are scanned to evaluate residuals $\tau_{j}=\min_{\mathbf{v}_{i}}\overline{\operatorname{res}}(z_{j},\Psi\mathbf{v}_{i})$.
+Comparatively, ResDMD’s lower cost ($O(N_{K}^{2}m+n_{z}N_{K}^{3})$) avoids iteration, but sacrifices dictionary adaptability.
+To mitigate these costs, we will consider some methods: (1) We consider Mini-Batch SGD since it uses a batch size $b\ll m$ that reduces per-iteration complexity to $O(N_{K}^{2}b+N_{K}^{3}+|\theta|)$, potentially lowering overall cost despite increased $k$, as validated by Barron space convergence properties.
+(2) We can use preconditioning to improve the condition number of $G(\theta)$ via diagonal scaling or incomplete Cholesky factorization could accelerate convergence, reducing $k$.
+Theoretical Extensions: The $O(1/n)$ SGD rate assumes convexity, yet $J(\theta)$’s non-convexity suggests sublinear convergence.
+A bound like $\|\mathcal{K}-\mathcal{K}_{N_{K}}\|$ (Appendix A.4) could quantify accuracy, while determining minimal $N_{K}$ and $k$ for spectral completeness remains a future work.
+Extending ResKoopNet to stochastic systems (e.g., adding noise terms to spectral residual) or embedding physical constraints (e.g., unitary spectra for Hamiltonian systems) could enhance its scope, respectively, which also aligns with future work outlined in the Section 5.
+A.7 Extra tests for ResDMD and ResKoopNet
+A.7.1 ResDMD with less dictionary functions
+While the original ResDMD implementation used 964 basis functions to capture the full spectrum of the nonlinear pendulum, our additional tests demonstrate that the unit circle spectrum can be reasonably approximated by ResDMD with fewer observables.
+Figure 7 shows the progressive improvement in spectral approximation as the dictionary size $N_{K}$ increases.
+With nearly 500 basis functions, ResDMD begins to adequately reconstruct the unit circle structure, though still requiring substantially more observables than the 300 basis functions needed by ResKoopNet.
+Refer to caption Refer to caption Refer to caption Refer to caption
+(a) $N_{K}=252$ (b) $N_{K}=356$ (c) $N_{K}=464$ (d) $N_{K}=860$
+Figure 7: Comparison across different dictionary sizes by ResDMD.
+A.7.2 ResKoopNet with different hyperparameters
+We test ResKoopNet on the dataset which has 90 initial points, as used in the Section 4.1.
+The Figure 8 shows that when tunning the number of hidden layers from 1 to 4 and the number of neurons in each layer from 250 to 350, the results change significantly.
+We can also tell that a suitable NN structure would be about 3 hidden layers with 300 neurons in each hidden layer.
+More neurons and hidden layers would be fore help, but cost more computational resource.
+This result also shows that our method is robust to hyperparameters tuning.
+Refer to caption 1 hidden layer 250 neuron Refer to caption 1 hidden layer 275 neuron Refer to caption 1 hidden layer 300 neuron Refer to caption 1 hidden layer 325 neuron Refer to caption 1 hidden layer 350 neuron
+Refer to caption 2 hidden layers 250 neuron Refer to caption 2 hidden layers 275 neuron Refer to caption 2 hidden layers 300 neuron Refer to caption 2 hidden layers 325 neuron Refer to caption 2 hidden layers 350 neuron
+Refer to caption 3 hidden layers 250 neuron Refer to caption 3 hidden layers 275 neuron Refer to caption 3 hidden layers 300 neuron Refer to caption 3 hidden layers 325 neuron Refer to caption 3 hidden layers 350 neuron
+Refer to caption 4 hidden layers 250 neuron Refer to caption 4 hidden layers 275 neuron Refer to caption 4 hidden layers 300 neuron Refer to caption 4 hidden layers 325 neuron Refer to caption 4 hidden layers 350 neuron
+Figure 8: Comparisons of neural network architectures.
+Each row corresponds to the number of hidden layers (from 1 to 4), and each column shows the model with a different neuron count (250, 275, 300, 325, 350).
+A.8 Hankel-DMD
+A.8.1 Justification of using Hankel-DMD as comparison in all experiments
 Hankel-DMD operates by constructing a Hankel matrix from time-delayed measurements of the system state, based on Takens’ embedding theorem, which states that time-delayed coordinates can reconstruct the state space of dynamical systems.
@@
 Consequently, we also apply Hankel-DMD to the turbulence and neural dynamics experiments to evaluate its effectiveness in these representative high-dimensional settings.
-As results, although its performance rivals NN-ResDMD in the simple pendulum system by showing eigenvalues points near the unit circle and containing some polluted eigenvalues, which are close to the ground truth unit circle, we would like to emphasize that it capture the point spectrum and miss the full spectral information.
+As results, although its performance rivals ResKoopNet in the simple pendulum system by showing eigenvalues points near the unit circle and containing some polluted eigenvalues, which are close to the ground truth unit circle, we would like to emphasize that it capture the point spectrum and miss the full spectral information.
 When it comes to high-dimensional systems, it fails to capture key dynamics in higher-dimensional systems, as seen in the later experiment (Section 4.2 and Section 4.3).
-A.6.2 Application in Turbulence
-Here we present the Koopman modes computed by Hankel-DMD for comparison with the NN-ResDMD results.
-As shown in Figure 7, despite having small residuals, these modes fail to clearly capture the fundamental pressure field structure that was successfully identified by NN-ResDMD’s first Koopman mode (see Figure 5).
-This comparison demonstrates the superior ability of NN-ResDMD to extract physically meaningful patterns from complex fluid systems.
-Figure 7: The plots illustrate turbulence detection using the four Koopman modes computed by Hankel-DMD, which are ranked with their corresponding residuals from the smallest.
-A.7 Practical details for neural data analysis
-A.7.1 Dataset Details and Experimental Setup
+A.8.2 Application in turbulence
+Here we present the Koopman modes computed by Hankel-DMD for comparison with the ResKoopNet results.
+As shown in Figure 9, despite having small residuals, these modes fail to clearly capture the fundamental pressure field structure that was successfully identified by ResKoopNet’s first Koopman mode (see Figure 5).
+The discrepancy observed for Hankel-DMD [1] is due to its fundamentally different theoretical framework.
+While ResKoopNet (and other EDMD-based methods) use a Galerkin framework with convergence guarantees established via minimizing the spectral residual, the Hankel-DMD approach is based on Takens’ embedding theorem [36].
+Therefore, the spectral residual metric, which is derived and justified within the Galerkin projection framework, does not fully extend to the Hankel-DMD setting.
+Even if Hankel-DMD yields small residual, this does not ensure that its modes accurately capture the underlying dynamics (such as the fundamental pressure field structure).
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 9: The plots illustrate turbulence detection using the four Koopman modes computed by Hankel-DMD, which are ranked with their corresponding spectral residual from the smallest, similar as in ResKoopNet case.
+A.9 Practical details for neural data analysis
+A.9.1 Dataset details and experimental setup
 The dataset utilized in this study is part of the open dataset provided for the ’Sensorium 2023’ competition [39].
@@
 The objective of the competition is to predict large-scale neuronal population activity in response to different frames of the stimulus videos, based on the hypothesis that population dynamics in the primary visual cortex, driven by visual stimuli, encode significant information about the dynamics of the videos [3, 29, 15].
-A.7.2 Task Definition and Rationale
+A.9.2 Task definition and rationale
 In contrast to the competition’s prediction objective, our study focuses on the task of state partitioning of neural signals.
-While prediction remains feasible, we aim to demonstrate that state partitioning is sufficient to highlight the superiority of NN-ResDMD over a series of other methods in uncovering the latent dynamics of the system.
+While prediction remains feasible, we aim to demonstrate that state partitioning is sufficient to highlight the superiority of ResKoopNet over a series of other methods in uncovering the latent dynamics of the system.
 Specifically, in each experiment, a set of six video stimuli was repeatedly presented to each mouse, creating ideal conditions for defining brain states.
@@
 This provides a reliable basis for testing the efficacy of Koopman decomposition methods in uncovering latent dynamics and distinguishing these states.
-A.7.3 Dataset Structure and Dimensionality
+A.9.3 Dataset structure and dimensionality
 The dataset includes neural recordings from five mice, with each mouse responding to six distinct video stimuli, presented in 9-10 repeated trials (resulting in approximately 60 trials in total).
@@
 Thus, the data to be analyzed consists of a high-dimensional time series with 7000+ observables per snapshot.
-A.7.4 Implementations of NN-ResDMD and other classical methods
-We compare here four methods: the proposed NN-ResDMD and three classical Koopman decomposition methods for high-dimensional systems: the Hankel-DMD, the EDMD with RBF basis, and the Kernel ResDMD.
+A.9.4 Implementations of ResKoopNet and other classical methods
+We compare here four methods: the proposed ResKoopNet and three classical Koopman decomposition methods for high-dimensional systems: the Hankel-DMD, the EDMD with RBF basis, and the Kernel ResDMD.
 We applied them to the 5 datasets, although with slightly different implementations and different dimensions of approximated Koopman invariance subspace.
-For NN-ResDMD, we train the dictionaries with all the snapshots recorded in each mouse such that the total snapshot number is the product of the snapshot number in one trial and the number of all trials.
+For ResKoopNet, we train the dictionaries with all the snapshots recorded in each mouse such that the total snapshot number is the product of the snapshot number in one trial and the number of all trials.
 This is to avoid overfitting with the small snapshot numbers within a trial.
 The high-dimensional data is first reduced to 300 dimensions with Singular Value Decomposition.
-The dimension of the Koopman subspace is chosen to be 601, consisting of 300 trained bases and 301 pre-chosen ones (constant and the first-degree polynomials of the SVD-ed 300 dimensions).
-The first 501 eigenfunctions sorted by the modulus of eigenvalues are selected to avoid spurious eigenvalues estimation due to noise.
+The dimension of the Koopman subspace is chosen to be 50 (to be compared with Hankel DMD fairly), consisting of 25 trained bases and 25 pre-chosen ones (constant and the first-degree polynomials of the SVD-ed 24 dimensions).
 One can find the decomposed eigenfunctions in Figure 6A(top), with a marker of the ground truth state separations based on stimulus identity.
 For Hankel-DMD, the Koopman eigenfunctions were approximated using the eigenvectors of the Hankel matrix.
-Specifically, the Hankel matrix was formed as in Equation 53 from arbabi2017ergodic [25], using all the observables from one trial of each mouse with a delay of 50.
+Specifically, the Hankel matrix was formed as in Equation 53 from Arbabi and Mezic [1], using all the observables from one trial of each mouse with a delay of 50.
 Consequently, the snapshot size became 249 times the observable number, and the resulting number of eigenfunctions was 50, each with a length of 50.
@@
 We chose the kernel function as the commonly-used normalized Gaussian function in the calculation.
-The Koopman eigenfunctions from both NN-ResDMD and other methods represent dynamical features corresponding to one of the six video stimuli.
+The Koopman eigenfunctions from both ResKoopNet and other methods represent dynamical features corresponding to one of the six video stimuli.
 To evaluate how well the eigenfunctions capture the latent dynamics, we assess the similarity of the features for trials with the same stimulus and their dissimilarity from those corresponding to different stimuli.
 Effectively, this makes the problem a clustering task, where the separability of the Koopman eigenfunctions reflects how well they capture the key dynamic components related to the stimuli.
-A.8 Choice justification of dictionary sizes
+A.10 Choice justification of dictionary sizes
 In this section, we provide justifications for the use of different dictionary sizes (i.e., the number of Koopman eigenfunctions) in the aforementioned four methods for the neural dynamics experiment.
@@
   1. 1.
-    For NN-ResDMD, we selected 300 trained basis functions and 300 first-order monomial basis functions as the dictionary for the 300 reduced observables.
-This choice ensures the dictionary is rich enough to span the Koopman invariant subspace.
-Hence, the size of the trained dictionary was set to be at least equal to the original observable size.
-Then based on the rank of estimated Koopman eigenvalues, we select the dominant 501 eigenfunctions to avoid the eigenfunctions with zero eigenvalues.
+    For ResKoopNet, we selected 25 trained basis functions and 24 first-order monomial basis functions as the dictionary for the 24 reduced observables.
+This choice ensures the dictionary size is the same as Hankel DMD, making it a fair comparison between the performance of these two methods.
+The last two methods uses higher dictionary sizes but still remain ineffective in capturing temporal dynamics, thus not affecting the fairness of choosing 50 dictionaries here.
+The size of the trained dictionary was set to be at least equal to the original observable size to ensure a sufficiently rich dictionary for the Koopman invariant subspace.
   2. 2.
     For Hankel DMD, the number of delays (as dictionary size/number of eigenfunctions) is first constrained by the temporal sample size (i.e., snapshot size) because it cannot exceed the maximum snapshot size.
-Therefore, it is impossible to choose the same dictionary size as the NN-ResDMD example.
+Therefore, it is impossible to choose the same dictionary size as the ResKoopNet example.
 Choosing the delay too small will result in an insufficient dictionary size to span the Koopman invariant subspace, and too large will reduce the actual snapshot size to estimate the covariance matrices in the estimation of the Koopman matrix.
@@
 However, our previous experience with a similar dataset and the results of using the RBF basis for the EDMD method all suggest that the performance will be better with more dictionary functions.
-Therefore, we chose 1000 RBF basis and the original 300 first-order monomial basis as a better condition compared to the same dictionary size with NN-ResDMD.
+Therefore, we chose 1000 RBF basis and the original 300 first-order monomial basis as a better condition compared to the same dictionary size with ResKoopNet.
   4. 4.
-    For Kernel ResDMD, the dictionary size is theoretically determined to be the number of snapshots. Therefore, we cannot make the dictionary size consistent with the NN-ResDMD example.
+    For Kernel ResDMD, the dictionary size is theoretically determined to be the number of snapshots. Therefore, we cannot make the dictionary size consistent with the ResKoopNet example.
 Based on the above justifications, we believe our choices of dictionary sizes are reasonable and ensure a fair comparison across the methods.
-A.8.1 Visualization and Clustering Performance
+A.10.1 Visualization and clustering performance
 To visualize the clustering of high-dimensional Koopman eigenfunctions, we perform dimensionality reduction using Multi-dimensional Scaling (MDS).
@@
 For illustration, we present the results of reducing the feature space to two dimensions (Figure 6B-E).
-The NN-ResDMD reduced features for the six types of trials (corresponding to the six video stimuli) are well-separated for all five mice (Figure 6B).
+The ResKoopNet reduced features for the six types of trials (corresponding to the six video stimuli) are well-separated for all five mice (Figure 6B).
 In contrast, the Hankel-DMD features show no clear clustering structure (Figure 6C).
 Similarly, the features produced by EDMD with an RBF basis and Kernel ResDMD do not show clear separability (Figure 6D-E, Appendix Figure 14B-D, Appendix Figure 15B-D).
-Figure 8: State Partition performance of eigenfunctions for NN-ResDMD and Hankel-DMD in 2D space visualized with UMAP (A) and t-SNE (B).
+Figure 10: Trial-averaged neural responses to 6 video stimuli (marked by 6 states).
+Figure 11: (A) Mean firing rates (calcium spike) of each mouse in response to 6 video stimuli (marked as states).
+(B) Clustering performance based on the mean firing rates of each mouse to 6 video stimuli, visualized with Multi-dimensional Scaling (MDS).
+(C) Same as (B) but visualized with UMAP.
+(D) Same as (B) but visualized with t-SNE
+Figure 12: Comparison of clustering quality based on Davies-Bouldin Indices for clustering with mean firing rates and Koopman eigenfunctions learned from ResKoopNet, visualized in 2D space with Multi-dimensional Scaling (A), UMAP (B) and t-SNE (C).
+Figure 13: State Partition performance with eigenfunctions for ResKoopNet (50 bases) and Hankel-DMD in 2D space visualized with UMAP (A) and t-SNE (B).
 Figure 14: Full results of EDMD with RBF basis.
@@
 Figure 15: Same as Figure 14 but estimated with Kernel ResDMD, with 299 basis of the Koopman subspace, thus 299 eigenfunctions.
-Figure 11: Davies-Bouldin Indices evaluating the clustering performance of dynamical components learned by four methods (NN-ResDMD, Hankel DMD, EDMD+RBF, and Kernel ResDMD) across five mice.
+Figure 16: Davies-Bouldin Indices evaluating the clustering performance of dynamical components learned by four methods (ResKoopNet(50 bases), Hankel DMD, EDMD+RBF, and Kernel ResDMD) across five mice.
 Comparisons are shown using UMAP (A) and t-SNE (B).
-A.8.2 Clustering Quality Metrics
+Figure 17: Davies-Bouldin Indices for different neural network hyperparameters in training ResKoopNet with 50 bases in an example mouse.
+Note that lower values indicate better clustering performance.
+With small layer size and layer number, the clustering performance is not stable but becomes robust when the layer size reached 200 and layer number reaches 3.
+A.10.2 Clustering quality metrics
 We further quantified the clustering quality by calculating the Davies-Bouldin Index (DBI) for both Koopman decomposition methods across all mice (Figure 6F).
@@
 A lower DBI indicates better clustering performance.
-NN-ResDMD features yield significantly lower DBI scores compared to other methods, confirming that NN-ResDMD produces more clearly defined clusters corresponding to the ground truth trials.
-Similar clustering results are observed with UMAP and t-SNE (see Appendix Figure 11), further supporting the superior performance of NN-ResDMD in capturing the latent dynamic structure compared to the other classical methods.
-A.9 Source code
-For reproducibility, the source code will be available at the following anonymous URL: https://anonymous.4open.science/r/ICLR-7305-PROJ.
-A full version of the codebase will be released upon acceptance of the paper.
+ResKoopNet features yield significantly lower DBI scores compared to other methods, confirming that ResKoopNet produces more clearly defined clusters corresponding to the ground truth trials.
+Similar clustering results are observed with UMAP and t-SNE (see Appendix Figure 16), further supporting the superior performance of ResKoopNet in capturing the latent dynamic structure compared to the other classical methods.

### v4

@@
 Abstract
-Analyzing long-term behaviors in high-dimensional nonlinear dynamical systems remains challenging, with the Koopman operator framework providing a powerful global linearization approach, though existing methods for approximating its spectral components often suffer from theoretical limitations and reliance on predefined dictionaries.
-While Residual Dynamic Mode Decomposition (ResDMD) introduced the spectral residual to assess the accuracy of Koopman operator approximation, it only filters precomputed spectra, which prevents it from fully discovering the Koopman operator’s complete spectral information (a limitation sometimes referred to as the ‘spectral inclusion’ problem).
-We introduce ResKoopNet (Residual-based Koopman-learning Network), a novel method that addresses this limitation by explicitly minimizing the spectral residual to compute Koopman eigenpairs, which can identify a more precise and complete spectrum of the Koopman operator.
-This approach provides theoretical guarantees while maintaining computational adaptability through a neural network implementation.
-Experiments on physical and biological systems demonstrate ResKoopNet’s superior accuracy in spectral approximation compared to existing methods, particularly for systems with continuous spectra and high dimensional, which makes it as an effective tool for analyzing complex dynamical systems.
+Analyzing the long-term behavior of high-dimensional nonlinear dynamical systems remains a significant challenge.
+While the Koopman operator framework provides a powerful global linearization tool, current methods for approximating its spectral components often face theoretical limitations and depend on predefined dictionaries.
+Residual Dynamic Mode Decomposition (ResDMD) advanced the field by introducing the spectral residual to assess Koopman operator approximation accuracy; however, its approach of only filtering precomputed spectra prevents the discovery of the operator’s complete spectral information, a limitation known as the ‘spectral inclusion’ problem.
+We introduce ResKoopNet (Residual-based Koopman-learning Network), a novel method that directly addresses this by explicitly minimizing the spectral residual to compute Koopman eigenpairs.
+This enables the identification of a more precise and complete Koopman operator spectrum.
+Using neural networks, our approach provides theoretical guarantees while maintaining computational adaptability.
+Experiments on a variety of physical and biological systems show that ResKoopNet achieves more accurate spectral approximations than existing methods, particularly for high-dimensional systems and those with continuous spectra, which demonstrates its effectiveness as a tool for analyzing complex dynamical systems.
 1 Introduction
@@
 As a result, EDMD can miss critical information about the underlying dynamics, particularly in systems exhibiting rich, complex behavior.
-To address these limitation, Residual Dynamic Mode Decomposition (ResDMD) [8] was introduced to offer convergence guarantees through a spectral residual that measures how well the estimated Koopman spectra are.
+To address these limitations, Residual Dynamic Mode Decomposition (ResDMD) [8] was introduced to offer convergence guarantees through a spectral residual that measures how well the estimated Koopman spectra are.
 By assessing convergence of spectral residual, ResDMD eliminates spurious spectral components that do not represent the system’s true dynamics, which enhances the reliability of spectral estimation.
@@
 […]
+$=\sum_{i,j=1}^{N_{K}}\bar{v}_{i}[\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\lambda\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\bar{\lambda}\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}+|\lambda|^{2}\langle\psi_{i},\psi_{j}\rangle_{\mu}]v_{j},$ (3.1)
 where $\bar{v}_{i},\bar{\lambda}$ denote the complex conjugates of $v_{i},\lambda$.
@@
 where * denotes conjugate transpose of matrix or adjoint operator. Using these relationships, the spectral residual in Eq. (3.1) can be approximated as (see A.2 for more details):
-$\begin{split}\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}\\ -\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}]\mathbf{v}.\end{split}$ (3.3)
+$\begin{split}\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}-\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}]\mathbf{v}.\end{split}$ (3.3)
 Notice that Eq.
@@
 The overall pipeline is illustrated in Figure 1.
-ResKoopNet optimizes the dictionary functions by minimizing the total spectral residual,
+ResKoopNet optimizes the dictionary functions by minimizing the total spectral residual $J$ as following
 $J\coloneqq\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2},$
@@
 Notice that the neural network architecture can be replaced by more advanced models or other optimizers depending on the system and data structure.
-Figure 1: The optimizer on the right looks for the optimal dictionary by minimizing Eq. (3.3).
+Figure 1: The optimizer on the right searches for the optimal dictionary by minimizing Eq. (3.3).
 In ResDMD’s framework, the spectral residual in Eq.
@@
 Output: $\tilde{K}(\theta)$, eigenpairs $\{(\lambda_{i},\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i})\}_{i=1}^{N_{K}}$ and pseudospectrum $\{z_{j}:\tau_{j}<\varepsilon\}$ (See Eq. (3.4)).
-Algorithm 1 NN-ResDMD
+Algorithm 1 ResKoopNet
 ResKoopNet’s practical advantages come with computational demands due to its iterative optimization.
@@
 Note that we include Hankel-DMD [1] as a consistent benchmark across all experiments due to its theoretical guarantees and applicability to both low and high-dimensional systems without requiring predefined dictionaries.
-It uses time-delayed state measurements by Takens Embedding Theorem [36] instead of using a pre-defined dictionary or applying Galerkin approximation as in EDMD and ResDMD methods (see Appendix A.8.1 for more details.)
+It uses time-delayed state measurements by Takens Embedding Theorem [35] instead of using a pre-defined dictionary or applying Galerkin approximation as in EDMD and ResDMD methods (see Appendix A.8.1 for more details).
 4.1 Pendulum
@@
 As shown in Figure 2, ResKoopNet requires only $N_{K}=300$ observables to approximate the full spectrum.
-While the original ResDMD work reported using 964 basis functions [8, Section 4.3.1], our additional tests indicate that approximately 460 basis functions can produce reasonable circular approximations with ResDMD (see Figure 7 in Appendix A.7.1).
+While the original ResDMD work reported using 964 observables [8, Section 4.3.1], our additional tests indicate that approximately 460 basis functions can produce reasonable circular approximations with ResDMD (see Figure 7 in Appendix A.7.1).
 Nevertheless, ResKoopNet still achieves comparable accuracy with significantly fewer observables.
@@
 Recovering spatial patterns is a typical goal of DMD-based methods, especially in fluid dynamics [32, 27].
-Figure 5(a) shows the ground truth pressure distribution, where a clear asymmetry exists between the upper and lower airfoil surfaces.
+Figure 5(a) shows the ground truth pressure distribution, in which a clear asymmetry exists between the upper and lower airfoil surfaces.
 This fundamental spatial pattern is critical for analyzing aerodynamic performance and structural integrity.
@@
 Lower DBI values for ResKoopNet indicate better clustering.
-Since ResKoopNet directly minimizes the residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics.
+Since ResKoopNet directly minimizes the spectral residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics.
 To evaluate how effectively ResKoopNet reveals latent temporal dynamics in real data, we apply it to a dataset of high-dimensional neural signals and demonstrate its advantages over a series of classical methods: the Hankel-DMD, EDMD (combined with RBF basis) and Kernel-ResDMD.
@@
 For ResKoopNet, we trained dictionaries on all snapshots from each mouse to avoid overfitting.
-We reduced the data to 24 dimensions via truncated SVD, together with the constant and 25 trainable bases, resulting in 50 eigenfunctions.
+We reduced the data to 24 dimension via truncated SVD, together with the constant and 25 trainable bases, resulting in 50 eigenfunctions.
 The neural network we use here consists of 3 hidden layers, each with 200 neurons.
-The hyperparameter scanning results are included in Appendix Figure 17, demonstrating that our parameter choice is within the robust range.
+The hyperparameter scanning results are included in Appendix Figure 17, which demonstrates that our parameter choice is within the robust range.
 The decomposed eigenfunctions are shown in Figure 6A(top), with markers indicating ground truth state separations.
@@
 These eigenfunctions, shown in Figure 6A(bottom), Appendix Figure 14A, and Appendix Figure 15A, are compared to ground truth trial identities.
-The Koopman eigenfunctions represent dynamical features corresponding to the video stimuli.
-To evaluate their effectiveness, we assess how well eigenfunctions of the same stimuli cluster together, which distinguishes them from other states.
-If the eigenfunctions capture key dynamics related to the stimuli, those from trials with the same video should be separable from others.
-This turns the problem into a clustering task based on the separability of eigenfunctions across different stimuli.
-Note that averaged trial differences are even visibly clear for the ResKoopNet case.
-It is worth clarifying that clustering these processes is not biologically necessary, as simple clustering of mean firing rates is sufficient (see Appendix Figures 1112).
+In this study, Koopman eigenfunctions are interpreted as representing dynamical features corresponding to the video stimuli.
+Consequently, their effectiveness in capturing these key, stimulus-related dynamics is evaluated by assessing how well eigenfunctions from trials with the same video stimulus cluster together and remain distinguishable from those associated with different video stimuli.
+This effectively transforms the problem into a clustering task based on the separability of eigenfunctions across different stimuli.
+It is noteworthy that for the ResKoopNet case, averaged trial differences are visibly clear.
+It is also worth clarifying that clustering these processes is not biologically necessary, as simple clustering of mean firing rates is sufficient (see Appendix Figures 101112).
 However, this dataset provides an ideal benchmark with clear ground truth labels, enabling accurate comparison of Koopman eigenfunction estimation methods.
-Effective estimation of Koopman eigenfunctions can be valuable for more complex tasks, such as unsupervised latent state identification or decoding object dynamics from video stimuli, offering promising directions for future research.
+Effective estimation of Koopman eigenfunctions can be valuable for more complex tasks, such as unsupervised latent state identification or decoding object dynamics from video stimuli, which offers promising directions for future research.
 We use Multi-dimensional Scaling (MDS) to visualize how these eigenfunction-based features cluster according to ground truth states.
@@
 In this paper, we introduced ResKoopNet, a novel approach that minimizes spectral residuals to accurately approximate Koopman operators for complex dynamical systems.
-This method successfully resolves the spectral inclusion problem while capturing both discrete and continuous spectra with high fidelity.
+This method successfully resolves the spectral inclusion problem while capturing both discrete and continuous spectra with high accuracy.
 Our experiments across physical and biological systems demonstrate that ResKoopNet significantly outperforms existing methods in identifying coherent structures and uncovering latent dynamics, particularly in high-dimensional applications.
-Despite these strengths, ResKoopNet has several limitations.
-First, the use of neural networks increases computational cost compared to classical algorithms (see Appendix A.6).
-Second, unlike stochastic approaches such as VAMP [25] and SDMD [46], ResKoopNet currently does not account for stochasticity (see Appendix A.5 for comparison).
+Despite these advantages, ResKoopNet has several limitations.
+First, the use of neural networks increases computational cost compared to classical numerical algorithms (see Appendix A.6).
+Second, unlike stochastic approaches such as VAMP [24] and SDMD [46], ResKoopNet currently does not account for stochasticity (see Appendix A.5 for comparison discussion).
 In addition, its performance can vary depending on the neural architecture and training configuration.
-Many physical laws, such as conservation principles or symmetry constraints, are inherently reflected in the spectral properties of dynamical systems.
-Incorporating such physical structure into the neural network architecture may therefore enhance the learning of Koopman spectral information.
-Future work could explore physics-informed designs that embed prior knowledge directly into the learning process, allowing the network to better align with the system’s underlying dynamics.
-This integration may improve both the interpretability and accuracy of the extracted Koopman representations.
+There are several promising directions to address these limitations and extend ResKoopNet’s capabilities.
+Many physical laws, such as conservation principles or symmetry constraints, are inherently encoded in the spectral properties of dynamical systems.
+Incorporating such domain knowledge directly into the neural network architecture could significantly enhance the learning of Koopman spectral information.
+Future work could explore physics-informed neural network designs, develop efficient training strategies to reduce computational costs, and extend the framework to handle stochastic dynamics.
+Ultimately, ResKoopNet and its future variants have the potential to become powerful tools for understanding and predicting the behavior of complex dynamical systems across diverse scientific and engineering applications, and bridge the gap between data-driven methods and physical principles.
 Acknowledgments
@@
 A.1 Source Code
-For reproducibility, the source code will be available at the following link.
+For reproducibility, the source code will be available at this link.
 A.2 Calculation steps for Eq. (3.3)
@@
 To solve $\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)$ , we simply rewrite it as
-$\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)=\tfrac{d}{dK}\operatorname{tr}\left((K\mathbf{v}_{i})^{*}G(K\mathbf{v}_{i})\right)$
+$\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)=\tfrac{d}{dK}\operatorname{tr}\left((K\mathbf{v}_{i})^{*}G(K\mathbf{v}_{i})\right).$
 A.4 Discussion on convergence analysis
@@
   2. (b)
-    The loss $J(\theta)$ is Lipschitz continuous in $\theta$, and the neural network $\Psi(x;\theta)$ has bounded gradients, facilitating gradient-based optimization.
+    The loss $J(\theta)$ is strongly convex and Lipschitz continuous in $\theta$, and the neural network $\Psi(x;\theta)$ has bounded gradients, facilitating gradient-based optimization.
 Now, consider a Barron space $\mathcal{B}$ which is dense in $\mathcal{F}$.
@@
 A.9.1 Dataset details and experimental setup
-The dataset utilized in this study is part of the open dataset provided for the ’Sensorium 2023’ competition [39].
+The dataset utilized in this study is part of the open dataset provided for the ‘Sensorium 2023’ competition [39].
 The dataset consists of calcium imaging recordings from the primary visual cortex of mice.